### v1

# June, 2024

June, 2024
Covariant quantisation of tensor multiplet models
Sergei M. Kuzenko and Emmanouil S. N. Raptakis
Department of Physics M013, The University of Western Australia
35 Stirling Highway, Perth W.A. 6009, Australia
Email: sergei.kuzenko@uwa.edu.au, emmanouil.raptakis@uwa.edu.au
The Batalin-Vilkovisky formalism is applied to quantise the ${\cal N}=1$ supersymmetric generalisation of the Freedman-Townsend (FT) model proposed by Lindström and Roček in 1983. This super FT theory describes a non-Abelian tensor multiplet and is known to be classically equivalent to a supersymmetric nonlinear sigma model. Using path integral considerations, we demonstrate that this equivalence holds at the quantum level. A modified Faddeev-Popov procedure is employed to quantise models for the ${\cal N}=2$ tensor multiplet in harmonic superspace.
1 Introduction
With the discovery of simple supergravity in four dimensions [1, 2] and its extended and higher-dimensional generalisations, new types of gauge theories (as compared with the standard Yang-Mills theories) have been introduced in quantum field theory. Their specific new features in the Lagrangian formalism are: (i) open gauge algebra; and/or (ii) linearly dependent gauge generators. These imply that covariant quantisation of such theories cannot be carried out using the standard Faddeev-Popov approach [3]. Methods for covariant quantisation of gauge theories with open gauge algebra were developed by Kallosh [4, 5] and de Wit and van Holten [6]. Quantisation of a general irreducible gauge with open algebra was accomplished by Batalin and Vilkovisky [7]. Reducible gauge theories (also known as gauge theories with linearly dependent generators) naturally arise when one deals with a gauge two-form in four dimensions [8, 9, 10] and, more generally, gauge $p$-forms in diverse dimensions. These naturally occur in supergravity theories, see, e.g., [11, 12, 13, 14] for early publications and [15, 16, 17] for reviews. Several consistent quantisation procedures have been developed to quantise reducible Abelian gauge theories such as gauge $p$-forms [18, 19, 20, 21, 22, 23], including the formulations proposed in [20, 23] which apply in the supersymmetric case.
The procedures developed in [18, 19, 20, 21, 22, 23] are not directly applicable for the quantisation of the model for a non-Abelian antisymmetric tensor gauge field proposed by Freedman and Townsend [24]. [^1] This work built on two earlier unpublished papers [25, 26]. An important property of the Freedman-Townsend (FT) theory is its classical equivalence to the principal chiral nonlinear sigma model. The same model had also been formulated by Thierry-Mieg in 1980 [21], however his work was published only ten years later. [^2] One of us (SMK) is grateful to Jean Thierry-Mieg for sharing the story of his work [21]. Thierry-Mieg’s work also built on the unpublished papers by Freedman [25] and Townsend [26].
An ideal procedure to quantise the FT model is the Batalin-Vilkovisky (BV) formalism [27], which is universally applicable to general reducible gauge theories with open algebra. Following earlier attempts to quantise the FT model [28, 29] using generalised Faddeev-Popov and BRST techniques, the BV quantisation of the FT model was carried out in [30, 31, 32] (see [33] for a review). [^3] Ref. [30], completed in August 1986, was one of the earliest applications of the BV formalism. It was accepted for publication in Sov. J. Nucl. Phys. in 1987. Shortly before publication it was withdrawn, after the authors had been informed by a colleague that the same problem had already been solved elsewhere. Due to a limited access to the journals, at the time it was not possible to verify this information, which in fact turned out to be false. Quantum equivalence of the FT model and the principal chiral nonlinear $\sigma$-model was studied in [30, 31, 34, 35]. In the Hamiltonian approach, quantisation of the FT model and its quantum equivalence to the $\sigma$-model was studied by Frolov and Slavnov [36].
The ${\cal N}=1$ supersymmetric extension of the FT model was introduced by Lindström and Roček in 1983 [37]. [^4] Actually, the authors of [37] gave credit to Cremmer and Ferrara, and Gates and Siegel for this model. Further studies and generalisations of this theory have been given in [38, 39, 40]. This is a non-Abelian extension of the ${\cal N}=1$ tensor multiplet proposed by Siegel [41].
Covariant quantisation of the free tensor multiplet model in a supergravity background was carried out in [23, 42]. The latter theory is classically equivalent to a non-conformal scalar multiplet model [41]. Their quantum equivalence was established in [42] in the case of an on-shell supergravity background, and in [23] for an arbitrary supergravity background.
In this paper we make use of the Batalin-Vilkovisky procedure to quantise the ${\cal N}=1$ super FT model, which describes a non-Abelian tensor multiplet. At the classical level, this theory is known to be equivalent to a supersymmetric nonlinear sigma model [38]. Using path integral considerations, we demonstrate that this equivalence holds at the quantum level.
To the best of our knowledge, an ${\cal N}=2$ supersymmetric extension of the FT model is not known in the literature. In this paper, a modified Faddeev-Popov procedure is employed to quantise models for the (Abelian) ${\cal N}=2$ tensor multiplet in harmonic superspace.
This paper is organised as follows. Section 2 is devoted to: (i) carrying out BV quantisation of the ${\cal N}=1$ super FT model (2.1); and (ii) proving its quantum equivalence to the supersymmetric nonlinear $\sigma$-model (2.11). Section 3 is devoted to covariant quantisation of ${\cal N}=2$ tensor multiplet models. For this goal we we make use of the modified Faddeev-Popov procedure described in [23]. Concluding comments are given in section 4. Rudiments of harmonic superspace are collected in appendix A.
2 Supersymmetric Freedman-Townsend model
The ${\cal N}=1$ supersymmetric extension [37] of the FT model is described by the action [^5] Our superspace notation and conventions mostly correspond to the textbooks [43, 44]. We also make use of the notational shorthands ${\rm d}^{4|4}z:={\rm d}^{4}x{\rm d}^{2}\theta{\rm d}^{2}\bar{\theta}$, ${\rm d}^{4|2}z_{+}:={\rm d}^{4}x{\rm d}^{2}\theta$ and ${\rm d}^{4|2}z_{-}:={\rm d}^{4}x{\rm d}^{2}\bar{\theta}$ to denote the integration measures of the Minkowski superspace and its chiral and antichiral subspaces.
$S_{0}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,V^{2}~{},$ (2.1)
where $\chi_{\alpha}$, $W_{\alpha}$ and $V={V}^{\dagger}$ are Lie algebra-valued superfields [^6] We will say that $\Upsilon$ is a Lie algebra-valued superfield if $\Upsilon=\Upsilon^{I}T^{I}$, where $\Upsilon^{I}$ is a superfield and $T^{I}$ are the generators of a semi-simple Lie group $G$. The latter are normalised as $\text{tr}(T^{I}T^{J})=\delta^{IJ}$ and satisfy the commutation relations $[T^{I},T^{J}]={\rm i}f^{IJK}T^{K}$, where $f^{IJK}$ are the totally antisymmetric structure constants. satisfying:
¯$\bar{\nabla}_{\dot{\alpha}}\chi_{\alpha}=0~{},\qquad\bar{\nabla}_{{\dot{\alpha}}}W_{\alpha}=0~{},\qquad\nabla^{\alpha}W_{\alpha}=\bar{\nabla}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}~{}.$ (2.2)
Operators $\nabla_{A}=(\nabla_{a},\nabla_{\alpha},\bar{\nabla}^{\dot{\alpha}})$ above are gauge-covariant derivatives and obey the algebra:
$\{\nabla_{\alpha},\nabla_{\beta}\}=0~{},\quad$$\{\nabla_{\alpha},\bar{\nabla}_{\dot{\alpha}}\}=-2{\rm i}\nabla_{\alpha{\dot{\alpha}}}~{},\quad\{\bar{\nabla}_{\dot{\alpha}},\bar{\nabla}_{\dot{\beta}}\}=0~{},$ (2.3a)
$[\nabla_{\alpha},\nabla_{\beta{\dot{\beta}}}]$ $={\rm i}\varepsilon_{\alpha\beta}\bar{W}_{\dot{\alpha}}~{},\qquad[\bar{\nabla}_{\dot{\alpha}},\nabla_{\beta{\dot{\beta}}}]={\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}W_{\alpha}~{},$ (2.3b)
$[\nabla_{\alpha{\dot{\alpha}}},\nabla_{\beta{\dot{\beta}}}]$ $=-\frac{1}{2}\varepsilon_{\alpha\beta}\bar{\nabla}_{({\dot{\alpha}}}\bar{W}_{{\dot{\beta}})}-\frac{1}{2}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\nabla_{(\alpha}W_{\beta)}~{}.$ (2.3c)
These constraints may be solved [45] in terms of the complex unconstrained prepotential $\Omega=\Omega^{I}T^{I}$ and its Hermitian conjugate $\bar{\Omega}\equiv\Omega^{\dagger}$,
$\nabla_{\alpha}={\rm e}^{-\Omega}D_{\alpha}{\rm e}^{\Omega}~{},\qquad\bar{\nabla}_{\dot{\alpha}}={\rm e}^{\bar{\Omega}}\bar{D}_{\dot{\alpha}}{\rm e}^{-\bar{\Omega}}~{},\qquad{\rm e}^{V}={\rm e}^{\Omega}{\rm e}^{\bar{\Omega}}~{},$ (2.4)
where $D_{A}=(\partial_{a},D_{\alpha},\bar{D}^{\dot{\alpha}})$ are the covariant derivatives of Minkowski superspace. [^7] Our normalisation of the gauge prepotentials in (2.4) differs from [44] and agrees with [46]. The superfields $\bar{\chi}_{\dot{\alpha}}$ and $\bar{W}_{\dot{\alpha}}$ in (2.1) are Hermitian conjugates of $\chi_{\alpha}$ and $W_{\alpha}$, respectively, in the vector representation. The conjugation rules are somewhat different in the chiral representation [46].
Making use of (2.2), it follows that action (2.1) is invariant under the following gauge transformation
$\delta_{L}\chi_{\alpha}=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}L~{},\qquad L=\bar{L}~{},$ (2.5)
where the gauge parameter $L=L^{I}T^{I}$ is Hermitian but otherwise unconstrained. It follows that the algebra of gauge transformations is Abelian and closed, $\delta_{L^{\prime}}\delta_{L}\chi_{\alpha}=0$. Actually, the super FT model is a reducible first-stage gauge theory, following the terminology of [27]. Indeed, shifting the gauge parameter in (2.5) as
$L~{}\to~{}L+\delta_{\sigma}L~{},\qquad\delta_{\sigma}L=\sigma+\bar{\sigma}~{},\quad\bar{\nabla}_{\dot{\alpha}}\sigma=0~{},$ (2.6a)
we obtain
$\delta_{\sigma}\delta_{L}\chi_{\alpha}={\rm i}\big{[}W_{\alpha},\sigma\big{]}~{},\qquad W_{\alpha}=-2\frac{\delta_{r}S_{0}}{\delta\chi^{\alpha}}~{},$ (2.6b)
and therefore $\delta_{\sigma}\delta_{L}\chi_{\alpha}$ vanishes on the mass shell. An important implication of (2.6b) is that the modified Faddeev-Popov quantisation procedures developed, e.g., in [20, 22, 23] cannot be directly applied to quantise the super FT model.
2.1 Classical equivalence
The tensor multiplet is known to admit self-interactions of the form [41, 37]
$S=\int{\rm d}^{4|4}z\,{\mathfrak{F}}(G)~{},\qquad G=\frac{1}{2}(D^{\alpha}\eta_{\alpha}+\bar{D}_{\dot{\alpha}}\bar{\eta}^{\dot{\alpha}})~{},\qquad\bar{D}_{\dot{\beta}}\eta_{\alpha}=0~{}.$ (2.7)
Here $\eta_{\alpha}$ denotes the chiral prepotential of a single tensor multiplet, to distinguish this case from the super FT model. The choice ${\mathfrak{F}}(G)=-\frac{1}{2}G^{2}$ corresponds to the free tensor multiplet, while ${\mathfrak{F}}(G)=-\mu\,G\log(G/\mu)$, with $\mu$ a positive constant, describes the so-called improved tensor multiplet [48]. The model has a classically equivalent (or dual) formulation [41, 37] realised in terms of a chiral scalar $\varphi$ and its conjugate $\bar{\varphi}$ with the action
$S=\int{\rm d}^{4|4}z\,{\mathfrak{U}}(\varphi+\bar{\varphi})~{},\qquad\bar{D}_{\dot{\beta}}\varphi=0~{},$ (2.8)
where ${\mathfrak{U}}$ is the Legendre transform of ${\mathfrak{F}}$ [37],
${\mathfrak{U}}={\mathfrak{F}}(h)-h(\varphi+\bar{\varphi})~{},\qquad{\mathfrak{F}}^{\prime}(h)=\varphi+\bar{\varphi}~{}.$ (2.9)
The target space of the supersymmetric $\sigma$-model (2.8) is a Kähler manifold with a $\mathsf{U}(1)$ isometry group consisting of holomorphic transformations $\varphi\to\varphi+{\rm i}r$, with $r\in{\mathbb{R}}$.
On general grounds [34], the models (2.7) and (2.8) should remain equivalent in the quantum theory. Direct studies of their quantum equivalence in a supergravity background was carried out in [42, 23] for the case ${\mathfrak{F}}(G)=-\frac{1}{2}G^{2}$.
It was shown in [38] that the super FT theory (2.1) is classically equivalent to a certain supersymmetric nonlinear sigma model. To see this, we first note that the dynamical equations for $\chi_{\alpha}$ and its conjugate imply that $V$ is pure gauge
$\frac{\delta_{l}S_{0}[\chi,\bar{\chi},V]}{\delta\chi^{\alpha}}=\frac{1}{2}W_{\alpha}=0\quad\implies\quad V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.10)
As a result, the action (2.1) reduces to
$S[\Phi,\bar{\Phi}]=\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,\log^{2}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (2.11)
This is a supersymmetric nonlinear $\sigma$-model with Kähler potential
$K(\Phi,\bar{\Phi})=\frac{1}{2}{\rm tr}\Big{(}\log^{2}({\rm e}^{\bar{\Phi}}{\rm e}^{{\Phi}})\Big{)}~{}.$ (2.12)
One of the goals of this section, which we pursue below, is to prove that this equivalence extends to the quantum level. Specifically, our goal is to demonstrate that the partition functions for models (2.1) and (2.11) coincide.
2.2 Quantisation via Batalin-Vilkovisky formalism
This subsection is devoted to quantising the super FT model (2.1). The covariant quantisation of tensor multiplet models (2.7) in a supergravity background was carried out in [42, 23] the modified Faddeev-Popov procedures. As $\delta_{\sigma}\delta_{L}\chi_{\alpha}$ vanishes only on the equations of motion, see eq. (2.6b), these quantisation procedures are no longer directly applicable in the super FT case. Thus, we will proceed by making use of the Batalin-Vilkovisky approach [27].
As a first step, it is necessary to mathematically describe the structure of this model’s gauge algebra. First, we recall that the generators of gauge transformations $R^{i}{}_{\underline{\alpha}}$ are defined via
$\delta_{\xi}\varphi^{i}=R^{i}{}_{\underline{\alpha}}\xi^{\underline{\alpha}}~{},$ (2.13)
where $\varphi=\{\chi^{\alpha I},\bar{\chi}_{\dot{\alpha}}^{I},V^{I}\}$ and $\xi=\{L^{I}\}$ denote the (classical) configuration space and gauge parameters, respectively. These generators can be read off from eq. (2.5):
$\delta_{L}\chi^{\alpha I}(z)$ $=\int{\rm d}^{4|4}z\,{R}^{\alpha I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}^{\alpha I\,J}(z,z^{\prime})=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla^{\alpha}\delta^{8}(z-z^{\prime})\delta^{IJ}~{},$ (2.14a)
$\delta_{L}\bar{\chi}_{\dot{\alpha}}^{I}(z)$ $=\int{\rm d}^{4|4}z\,{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})=\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\delta^{8}(z-z^{\prime})\delta^{IJ}~{}.$ (2.14b)
We also recall that, in the present model, the gauge parameters $L^{I}$ possess their own gauge symmetry, see eq. (2.6). This may be equivalently understood as the property that $R^{i}{}_{{\underline{\alpha}}}$ has nontrivial null eigenvectors $Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}$ at a stationary point of the action:
$\delta_{\zeta}\xi^{\underline{\alpha}}=Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}\zeta^{\underline{\alpha_{1}}}\quad\iff\quad R^{i}{}_{{\underline{\alpha}}}Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}=2\frac{\delta_{r}S_{0}}{\delta\varphi^{j}}K^{ji}_{\underline{\alpha}_{1}}(-1)^{\epsilon_{j}}~{},$ (2.15)
where the $K^{ji}_{\underline{\alpha}}$ term expresses the linear dependence of the generators off-shell. In the present case, the nontrivial null eigenvectors $Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}$ of the generators are:
$\delta_{\sigma}L^{I}(z)$ $=\int{\rm d}^{4|2}z_{+}\,Z^{I\,J}(z,z^{\prime})\sigma(z^{\prime})+\int{\rm d}^{4|2}z_{-}\,\bar{Z}^{I\,J}(z,z^{\prime})\bar{\sigma}(z^{\prime})~{},$ (2.16a)
$\implies$ $\quad Z^{I\,J}(z,z^{\prime})=\delta_{+}(z,z^{\prime})\delta^{IJ}~{},\qquad\bar{Z}^{I\,J}(z,z^{\prime})=\delta_{-}(z,z^{\prime})\delta^{IJ}~{}.$ (2.16b)
Finally, the $K^{ji}_{\underline{\alpha}_{1}}$ terms may be readily extracted from eq. (2.6):
$K^{\alpha I\,\beta J\,K}(z,z^{\prime},z^{\prime\prime})$ $=\varepsilon^{\alpha\beta}f^{IJK}\delta_{+}(z,z^{\prime})\delta_{+}(z,z^{\prime\prime})~{},$ (2.17a)
$\bar{K}_{{\dot{\alpha}}\,{\dot{\beta}}}^{I\,J\,K}(z,z^{\prime},z^{\prime\prime})$ $=-\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}f^{IJK}\delta_{-}(z,z^{\prime})\delta_{-}(z,z^{\prime\prime})~{}.$ (2.17b)
[…]
$\frac{\delta_{r}S}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\frac{\delta_{l}S}{\delta\mbox{\boldmath$\Phi$}_{\underline{A}}^{*}}=0~{}.$ (2.18)
Then, the in-out vacuum amplitude for $S_{0}[\varphi]$ may be defined as
$Z=\int\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{\underline{A}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}\big{]}\,\mu\,\delta\Big{[}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}-\frac{\delta\Psi}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\Big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]\Big{)}~{},$ (2.19)
where $\mu$ is the quantum integration measure and the fermionic functional $\Psi[\mbox{\boldmath$\Phi$}]$ is the gauge fermion, whose only requirement is that $Z$ possesses no residual gauge freedom.
For first-stage reducible theories a proper solution of the master equation (2.18) always exists for the so-called “minimal” set of fields
[…]
which, in the case of an Abelian and closed gauge algebra, takes the following form
$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]=S_{0}[\varphi]+\varphi^{*}_{i}R^{i}{}_{\underline{\alpha}}C^{\underline{\alpha}}+C^{*}_{\underline{\alpha}}Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}\eta^{\underline{\alpha}_{1}}+\varphi^{*}_{i}\varphi^{*}_{j}K^{ji}_{\underline{\alpha}_{1}}\eta^{\underline{\alpha}_{1}}~{}.$ (2.21)
Making use of equations (2.14), (2.16) and (2.17), we find $S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ to be:
$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ $=S_{0}[\chi,\bar{\chi},V]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}{\rm i}\big{(}\nabla^{\alpha}\chi_{\alpha}^{*}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{*{\dot{\alpha}}}\big{)}C+C^{*}(\eta+\bar{\eta})\Big{\}}$
$\phantom{=}+{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\chi^{*\alpha},\chi_{\alpha}^{*}\big{]}\eta-{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\bar{\chi}^{*}_{\dot{\alpha}},\bar{\chi}^{{\dot{\alpha}}*}\big{]}\bar{\eta}~{},$ (2.22)
where the ghost superfield $\eta$ is covariantly chiral, ¯$\bar{\nabla}_{\dot{\alpha}}\eta=0$.
[…]
$(i):$ $\qquad(\tilde{C},\pi)~{},\quad\epsilon(\tilde{C})=\epsilon(\pi)+1=1~{},$ (2.23a)
[…]
$(iii):$ $\qquad({\eta}^{\prime},{\pi}^{\prime})~{},\quad\bar{\nabla}_{\dot{\alpha}}\eta^{\prime}=0~{},\quad\bar{\nabla}_{\dot{\alpha}}\pi^{\prime}=0~{},\qquad\epsilon({\eta}^{\prime})=\epsilon({\pi}^{\prime})-1=0~{},$ (2.23c)
$(iv):$ $\qquad(\Lambda,\Pi)~{},\quad\bar{\nabla}_{\dot{\alpha}}\Lambda=0~{},\quad\bar{\nabla}_{\dot{\alpha}}\Pi=0~{},\qquad\epsilon(\Lambda)=\epsilon(\Pi)+1=1~{},$ (2.23d)
$(v):$ $\qquad(\Lambda^{\prime},\Pi^{\prime})~{},\quad\bar{\nabla}_{\dot{\alpha}}\Lambda^{\prime}=0~{},\quad\bar{\nabla}_{\dot{\alpha}}\Pi^{\prime}=0~{},\qquad\epsilon(\Lambda^{\prime})=\epsilon(\Pi^{\prime})+1=1~{}.$ (2.23e)
These will be used to obtain a new solution of master equation (2.18) that is more useful for the elimination of antifields. It is then easily seen that the following functional
$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]$ $=S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\tilde{C}^{*}\pi+\mbox{\boldmath$\eta$}^{*}\bar{\mbox{\boldmath$\pi$}}+\bar{\mbox{\boldmath$\eta$}}^{*}\mbox{\boldmath$\pi$}+\eta^{\prime*}\bar{\pi}^{\prime}+\bar{\eta}^{\prime*}\pi^{\prime}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\Lambda^{*}\bar{\Pi}+\bar{\Lambda}^{*}\Pi+\Lambda^{\prime*}\bar{\Pi}^{\prime}+\bar{\Lambda}^{*}\Pi^{\prime}\Big{\}}$ (2.24)
is also a solution to the master equation (2.18). It should be noted that this functional satisfies
$\Delta S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]=0~{},\qquad\Delta:=\frac{\delta_{r}}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\frac{\delta_{l}}{\delta{\mbox{\boldmath$\Phi$}}^{*}_{\underline{A}}}~{},$ (2.25)
therefore the quantum measure in (2.19) may be set to unity, $\mu=1$, in accordance with [47]. Choosing the gauge fermion to be
$\Psi(\mbox{\boldmath$\Phi$})$ $=\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{{\rm i}}{2}\tilde{C}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+C(\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}})+\tilde{C}(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\tilde{C}\pi\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\Big{\{}-\frac{1}{n}\big{(}{\mbox{\boldmath$\eta$}}\pi^{\prime}+\mbox{\boldmath$\pi$}\eta^{\prime}\big{)}+\Lambda\Pi+\Lambda^{\prime}\Pi^{\prime}\Big{\}}$
$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}-\frac{1}{n}\big{(}\bar{\mbox{\boldmath$\eta$}}\bar{\pi}^{\prime}+\bar{\mbox{\boldmath$\pi$}}\bar{\eta}^{\prime}\big{)}+\bar{\Lambda}\bar{\Pi}+\bar{\Lambda}^{\prime}\bar{\Pi}^{\prime}\Big{\}}~{},$ (2.26)
and eliminating antifields in eq. (2.2) leads to
$S[\mbox{\boldmath$\Phi$},{\mbox{\boldmath$\Phi$}}^{*}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta-\bar{\Pi}\Pi-\bar{\Pi}^{\prime}\Pi^{\prime}$
$\qquad\qquad\qquad+\Big{(}\frac{{\rm i}}{2}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\pi\Big{)}\pi$
$\qquad\qquad\qquad+\Big{(}-\frac{1}{2n}\pi^{\prime}-\frac{1}{4}\bar{\nabla}^{2}C\Big{)}\bar{\mbox{\boldmath$\pi$}}+\Big{(}-\frac{1}{2n}\bar{\pi}^{\prime}-\frac{1}{4}\nabla^{2}C\Big{)}\mbox{\boldmath$\pi$}$
$\qquad\qquad\qquad+\Big{(}\frac{1}{2n}\mbox{\boldmath$\pi$}-\frac{1}{4}\bar{\nabla}^{2}\tilde{C}\Big{)}\bar{\pi}^{\prime}+\Big{(}\frac{1}{2n}\bar{\mbox{\boldmath$\pi$}}-\frac{1}{4}\nabla^{2}\tilde{C}\Big{)}\pi^{\prime}\Big{\}}$
$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.27)
[…]
$Z$ $=\int\big{[}\mathscr{D}\mbox{\boldmath$\varphi$}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\varphi$}]\Big{)}~{},\qquad\mbox{\boldmath$\varphi$}=\Big{\{}\chi^{\alpha},\bar{\chi}_{\dot{\alpha}},V,C,\tilde{C},\eta,\bar{\eta},\mbox{\boldmath$\eta$},\bar{\mbox{\boldmath$\eta$}},\eta^{\prime},\bar{\eta}^{\prime}\Big{\}}~{},$
$S[\mbox{\boldmath$\varphi$}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}+\frac{n}{2}\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta$
$\qquad\qquad\qquad\qquad+m\bar{\eta}^{\prime}\eta^{\prime}-\frac{m}{8}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}^{2}\Big{\}}$
$\phantom{=}~{}-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.28)
[…]
2.3 Quantum equivalence
Having quantised the super FT model (2.1) above, we now specialise our analysis to the extension of classical equivalence demonstrated in section 2.1 to the quantum level. Our analysis will be similar to that given in [30] in the non-supersymmetric case.
[…]
$Z$ $=\int\big{[}\mathscr{D}V\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\big{[}\mathscr{D}\bar{\eta}\big{]}\big{[}\mathscr{D}\eta\big{]}\big{[}\mathscr{D}\bar{\mbox{\boldmath$\eta$}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\eta$}\big{]}\big{[}\mathscr{D}\pi\big{]}\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)\delta(\bar{W}_{\dot{\alpha}}-\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\pi)$
×$\qquad\qquad\times\delta(\nabla^{2}\pi)\delta(\bar{\nabla}^{2}\pi)\text{Det}(\mathcal{H})\exp\Big{(}\frac{{\rm i}}{\hbar}\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}V^{2}-2\tilde{C}\Box_{V}C+{\mbox{\boldmath$\eta$}}\bar{\eta}+{\eta}\bar{\mbox{\boldmath$\eta$}}\Big{\}}\Big{)}~{},$ (2.29)
where we have introduced the operators
$\mathcal{H}$ $=\begin{pmatrix}0&-\frac{1}{4}\bar{\nabla}^{2}\\ -\frac{1}{4}\nabla^{2}&0\end{pmatrix}~{},$ (2.30a)
$\Box_{V}$ $=-\frac{1}{16}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}-\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}$
$=\nabla^{a}\nabla_{a}-\frac{1}{4}W^{\alpha}\nabla_{\alpha}+\frac{1}{4}\bar{W}_{\dot{\alpha}}\bar{\nabla}^{\dot{\alpha}}~{}.$ (2.30b)
The operator $\mathcal{H}$ acts on the space of covariantly chiral-antichiral pairs in the adjoint representation of $G$, while $\Box_{V}$ is defined to act on the space of unconstrained scalar superfields in the adjoint representation of $G$.
To evaluate the $\delta$-functions in (2.3), we introduce a background quantum splitting [49] of the prepotentials in (2.4)
${\rm e}^{\Omega}={\rm e}^{\omega}\,{\rm e}^{\Omega_{Q}}~{},\qquad{\rm e}^{V}={\rm e}^{\omega}\,{\rm e}^{V_{Q}}\,{\rm e}^{\bar{\omega}}~{}.$ (2.31)
Here $\omega$ denotes the background prepotential and $V_{Q}$ the quantum gauge superfield. We denote by ${\cal D}_{A}=({\cal D}_{a},{\cal D}_{\alpha},\bar{\cal D}^{\dot{\alpha}})$ the background covariant derivatives. We will use the vector representation for the background covariant derivatives, and the chiral quantum representation for the original operators $\nabla_{A}$,
$\nabla_{\alpha}={\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}~{},\qquad\bar{\nabla}_{\dot{\alpha}}=\bar{\cal D}_{\dot{\alpha}}~{}.$ (2.32)
The field strength $W_{\alpha}$ takes the form
$W_{\alpha}={\rm e}^{\bar{\Omega}_{Q}}\left(-\frac{1}{4}\bar{\cal D}^{2}\Big{(}{\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}\Big{)}+w_{\alpha}\right){\rm e}^{-\bar{\Omega}_{Q}}~{},$ (2.33)
where $w_{\alpha}$ is the covariantly chiral background field strength, $\bar{\cal D}_{\dot{\beta}}w_{\alpha}=0$. Since we are interested in evaluating the $\delta$-functions in (2.3), we decompose $W_{\alpha}$ to first order in the quantum superfields,
$W_{\alpha}\approx-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+w_{\alpha}+\big{[}\bar{\Omega}_{Q},w_{\alpha}\big{]}~{}.$ (2.34)
The $\delta$-functions in (2.3) tell us that the background connection must be chosen to be flat,
$w_{\alpha}=0\quad\implies\quad{\rm e}^{\bar{\omega}}={\rm e}^{\Phi}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.35)
In evaluating $\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)$, it suffices to keep only the terms of first order in the quantum prepotentials and $\pi$,
$\delta\left(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi\right)=\delta\left(-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+\frac{{\rm i}}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}\pi\right)~{}.$ (2.36)
Now we consider the following change of variables:
$V_{Q}$ $=\rho+\bar{\rho}+\frac{1}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}+\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\rho=0~{},$ (2.37a)
$\pi$ $=\gamma+\bar{\gamma}+\frac{{\rm i}}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}-\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\gamma=0~{}.$ (2.37b)
Such a change of variables has a non-trivial Jacobian $\mathfrak{J}$. It can be computed by applying this change of variables in the right-hand side of
$1=\int\big{[}\mathscr{D}V_{Q}\big{]}\big{[}\mathscr{D}\pi\big{]}\exp\Big{(}\frac{{\rm i}}{2}\text{tr}\int\,{\rm d}^{4|4}z\,\Big{\{}{V_{Q}}^{2}-\pi^{2}\Big{\}}\Big{)}$ (2.38)
to result with
$\mathfrak{J}=\text{Det}(\widetilde{{\cal H}})\text{Det}_{+}^{-1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{-1/2}(\widetilde{\mathscr{H}}_{a})~{},$ (2.39)
where ~$\widetilde{\mathscr{H}}_{c}$ (~$\widetilde{\mathscr{H}}_{a}$) is a d’Alembertian on the space of covariantly chiral (antichiral) spinor superfields
~$\widetilde{\mathscr{H}}_{c}=\frac{1}{16}\bar{{\cal D}}^{2}{\cal D}^{2}~{},\qquad\widetilde{\mathscr{H}}_{a}=\frac{1}{16}{\cal D}^{2}\bar{{\cal D}}^{2}~{}.$ (2.40)
Here and below, a tilde over operators, such as $\widetilde{{\cal H}}$ and ~$\widetilde{{\square}}_{V}={\cal D}^{a}{\cal D}_{a}$, indicates that these operators are constructed using the flat-connection covariant derivatives ${\cal D}_{A}$. With this change of variables, we find that eq. (2.3) simplifies to
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{{\square}}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.41)
The quantum gauge field is now equal to zero, $V_{Q}=0$, and the relations (2.31) and (2.35) tell us that
$V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.42)
Now, repeating the analysis in [23] gives
$\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{\square}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})=1~{},$ (2.43)
and we obtain the in-out vacuum amplitude for the supersymmetric nonlinear $\sigma$-model (2.11)
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.44)
Thus we have demonstrated that the partition functions of the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.11) coincide, which implies the equality of the “on-shell” effective actions in these theories[34]. This establishes the quantum equivalence of the two theories. One may also add a source term for $V$ to (2.1) and repeat the analysis to obtain formal equality of their partition functions.
The analysis in this section may be generalised by coupling the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.11) to a background supergravity multiplet. Then (2.43) will turn into a topological invariant. As a result, the quantum supercurrents in the two theories $\langle J_{\alpha{\dot{\alpha}}}\rangle$ will coincide. This is an extension of the results obtained in [23, 42].
3 Quantisation of ${\cal N}=2$ tensor multiplet models
The ${\cal N}=2$ tensor multiplet [50] can be described by its gauge-invariant field strength $G^{ij}$ which is defined to be a real ${\mathsf{SU}}(2)$ triplet (that is, $G^{ij}=G^{ji}$ and ${\bar{G}}_{ij}:=\overline{G^{ij}}=G_{ij}$) subject to the covariant constraints [51, 52]
$D^{(i}_{\alpha}G^{jk)}={\bar{D}}^{(i}_{\dot{\alpha}}G^{jk)}=0~{},$ (3.1)
where $D_{A}=(\partial_{a},D_{\alpha}^{i},\bar{D}^{{\dot{\alpha}}}_{i})$ are covariant derivatives of ${\cal N}=2$ Minkowski superspace. These constraints are solved in terms of a chiral prepotential $\Psi$ [53, 54, 55, 56] as follows:
$G^{ij}=\frac{1}{8}D^{ij}\Psi+\frac{1}{8}D^{ij}{\bar{\Psi}}~{},\qquad{\bar{D}}^{i}_{\dot{\alpha}}\Psi=0~{},$ (3.2)
where we have used the following notation: $D^{ij}:=D^{\alpha(i}D_{\alpha}^{j)}$ and ${\bar{D}}^{ij}:=\bar{D}_{\dot{\alpha}}^{(i}\bar{D}^{{\dot{\alpha}}j)}$. The chiral prepotential $\Psi$ is defined modulo gauge transformations of the form
$\Psi\to\Psi+{\rm i}\Lambda~{},\qquad{\bar{D}}_{\dot{\alpha}}^{i}\Lambda=0~{},\qquad D^{ij}\Lambda={\bar{D}}^{ij}{\bar{\Lambda}}~{},$ (3.3)
which leave $G^{ij}$ invariant. Here the gauge parameter $\Lambda$ is a reduced chiral superfield, that is it satisfies the same constraints as the chiral field strength of the Abelian vector multiplet [57].
3.1 Is there an ${\cal N}=2$ supersymmetric FT model?
The algebra of ${\cal N}=2$ gauge-covariant derivatives $\nabla_{A}=(\nabla_{a},\nabla_{\alpha}^{i},\bar{\nabla}^{\dot{\alpha}}_{i})$ is [57]
$\{\nabla^{i}_{\alpha},\bar{\nabla}_{{\dot{\alpha}j}}\}=-2{\rm i}\delta^{i}_{j}\nabla_{\alpha{\dot{\alpha}}}~{},$ (3.4a)
$\{\nabla^{i}_{\alpha},\nabla^{j}_{\beta}\}={\rm i}\varepsilon_{\alpha\beta}\varepsilon^{ij}{\bar{W}}~{},\qquad\{{\bar{\nabla}}_{{\dot{\alpha}}i},{\bar{\nabla}}_{{\dot{\beta}}j}\}={\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\varepsilon_{ij}W~{},$ (3.4b)
${[}\nabla_{\alpha{\dot{\alpha}}},\nabla^{j}_{\beta}]=\frac{1}{2}\varepsilon_{\alpha\beta}{\bar{\nabla}}^{i}_{\dot{\alpha}}{\bar{W}}~{},\qquad[\nabla_{\alpha{\dot{\alpha}}},{\bar{\nabla}}_{{\dot{\beta}}i}]=\frac{1}{2}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\nabla_{\alpha i}W~{},$ (3.4c)
where $W$ is a covariantly chiral superfield strength, ¯$\bar{\nabla}^{\dot{\alpha}}_{i}W=0$, satisfying the Bianchi identity $\nabla^{ij}W=\bar{\nabla}^{ij}\bar{W}$. It follows from (3.4b) that a covariantly chiral matter multiplet $\phi$, transforming in some representation of the gauge group $G$, does not exist,
¯$\bar{\nabla}^{\dot{\alpha}}_{i}\phi=0\quad\implies\quad W\phi=0~{}.$ (3.5)
This means that there is no obvious ${\cal N}=2$ generalisation of the first two terms in (2.1). We therefore restrict our attention to Abelian tensor multiplets.
3.2 The free tensor multiplet model
The most general self-couplings of ${\cal N}=2$ tensor multiplets derived in [37] possess a simple manifestly ${\cal N}=2$ supersymmetric formulation within the so-called projective superspace approach [58, 59]. From the point of view of the quantum theory, the harmonic-superspace [60, 61] is advantageous. The two approaches are often complementary [62, 63].
Here we will use the harmonic-superspace formulation for the tensor multiplet [64, 61]. A free ${\cal N}=2$ tensor multiplet is described by the action
$S[\Psi,\bar{\Psi}]=\frac{1}{2}\int{\rm d}\zeta^{(-4)}\,(G^{++})^{2}~{},$ (3.6)
where the field strength (3.2) now takes the form
$G^{++}(z,u)=G^{ij}(z)\,u^{+}_{i}u^{+}_{j}=\frac{1}{8}(D^{+})^{2}\Psi(z)+\frac{1}{8}({\bar{D}}^{+})^{2}{\bar{\Psi}}(z)~{},$ (3.7a)
and the constraints (3.1) turn into
$D^{+}_{\alpha}G^{++}=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}G^{++}=0~{}.$ (3.7b)
It also follows from (3.7a) that
$D^{++}G^{++}=0~{}.$ (3.7c)
The action (3.6) can be written in a first-order form
$S[\Psi,\bar{\Psi},V^{++}]$ $=$ $\int{\rm d}\zeta^{(-4)}\,\left\{G^{++}V^{++}-\frac{1}{2}(V^{++})^{2}\right\}~{},$ (3.8)
where $V^{++}$ is a real analytic superfield of $\mathsf{U}(1)$ charge $+2$. The first term in this action can be rewritten as a chiral integral
$\int{\rm d}\zeta^{(-4)}\,G^{++}V^{++}=\frac{1}{2}\left\{\int{\rm d}^{4}x{\rm d}^{4}\theta\,\Psi W+{\rm c.c.}\right\}~{}.$ (3.9)
Here $W(z)$ is the (harmonic independent) chiral field strength of the ${\cal N}=2$ vector multiplet [57],
${\bar{D}}^{i}_{\dot{\alpha}}W=0~{},\qquad D^{ij}W={\bar{D}}^{ij}{\bar{W}}~{},$ (3.10)
which is expressed via the analytic prepotential $V^{++}$ as follows [60, 61, 65]:
$W(z)={1\over 4}\int{\rm d}u\,({\bar{D}}^{-})^{2}\,V^{++}(z,u)~{}.$ (3.11)
The gauge freedom (3.3) may be fixed by imposing the gauge condition $H^{++}=0$, where
$H^{++}=\frac{{\rm i}}{8}(D^{+})^{2}\Psi-\frac{{\rm i}}{8}({\bar{D}}^{+})^{2}\,{\bar{\Psi}}~{}.$ (3.12)
Choosing a gauge-fixing term
$S_{\rm gf}[\Psi,\bar{\Psi}]=\frac{1}{2}\int{\rm d}\zeta^{(-4)}\,(H^{++})^{2}~{},$ (3.13)
we obtain
$S[\Psi,\bar{\Psi}]+S_{\rm gf}[\Psi,\bar{\Psi}]=\frac{1}{2}\int{\rm d}^{4}x{\rm d}^{4}\theta{\rm d}^{4}\bar{\theta}\,\bar{\Psi}\Psi\equiv\frac{1}{2}\int{\rm d}^{4|8}z\,\bar{\Psi}\Psi~{}.$ (3.14)
Here we have used the integration rule
$\int{\rm d}^{4}x{\rm d}^{4}\theta{\rm d}^{4}\bar{\theta}\,\mathscr{L}=\int{\rm d}\zeta^{(-4)}\,(D^{+})^{4}\mathscr{L}~{},$ (3.15)
where the fourth-order operator $(D^{+})^{4}$ is defined in (A.6). Gauge-fixed action (3.14) is a higher-derivative model.
The tensor multiplet is known to be dual to a real hypermultiplet, see e.g. [37]. Within the harmonic-superspace approach, the tensor multiplet is dual to the so-called $\omega$ hypermultiplet, which was introduced in [60] and is described by the action
$S_{\omega}=-\frac{1}{2}\int{\rm d}\zeta^{(-4)}\,(D^{++}\omega)^{2}~{},\qquad D^{+}_{\alpha}\omega=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}\omega=0~{},$ (3.16)
where $\omega$ is subject to be real (with respect to the analyticity-preserving conjugation). The duality is manifested by considering the first-order action [64]
$S=\int{\rm d}\zeta^{(-4)}\,\left\{\frac{1}{2}(U^{++})^{2}+U^{++}D^{++}\omega\right\}~{},\qquad D^{+}_{\alpha}{U}^{++}=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}{U}^{++}=0~{},$ (3.17)
with $U$ a real analytic superfield.
3.3 Quantisation of the free tensor multiplet
The tensor multiplet model (3.6) is a gauge theory with linearly dependent generators. The gauge parameter $\Lambda$ in (3.3) is a reduced chiral superfield which can be expressed in terms of a real analytic superfield $\mathfrak{U}^{++}$ of $\mathsf{U}(1)$ charge $+2$,
$\Lambda[{\mathfrak{U}}^{++}]=\frac{1}{4}\int{\rm d}u\,({\bar{D}}^{-})^{2}{\mathfrak{U}}^{++}~{},\qquad D^{+}_{\alpha}{\mathfrak{U}}^{++}=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}{\mathfrak{U}}^{++}=0~{}.$ (3.18)
The chiral scalar $\Lambda$ does not change upon the replacement
${\mathfrak{U}}^{++}\to{\mathfrak{U}}^{++}+D^{++}\sigma~{},\qquad D^{+}_{\alpha}\sigma=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}\sigma=0~{}.$ (3.19)
This means that (3.6) is a first-stage reducible theory, and therefore it cannot be quantised using the Faddeev-Popov procedure [3]. In what follows, we will need another representation for the reduced chiral scalar (3.18), see e.g. [61]. It is
$\Lambda[{\mathfrak{U}}^{++}]=\frac{1}{4}({\bar{D}}^{+})^{2}{\mathfrak{U}}^{--}(z,u)~{},\qquad{\mathfrak{U}}^{--}(z,u)=\int{\rm d}u^{\prime}\,\frac{{\mathfrak{U}}^{++}(z,u^{\prime})}{(u^{+}u^{\prime+})^{2}}~{}.$ (3.20)
Several consistent quantisation procedures have been developed to quantise reducible Abelian gauge theories such as gauge $p$-forms [18, 19, 20, 21, 22, 23], including the formulations of [20, 23] which have been applied in the ${\cal N}=1$ supersymmetric case. These quantisation schemes are much easier to deal with than the Batalin-Vilkovisky formalism [27]. Here we will make use of the modified Faddeev-Popov procedure described in [23] to quantise the tensor multiplet model (3.6).
Our first step is to define a generalised delta-function $\hat{\delta}[H^{++}]$, with $H^{++}$ the gauge fixing function (3.12). The latter is a real analytic superfield obeying the constraint $D^{++}H^{++}=0$, and therefore the naive delta-function
$\delta\big{[}H^{++}\big{]}=\int\big{[}\mathscr{D}{\mathfrak{V}}^{++}\big{]}\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}H^{++}\right\}~{},\qquad D^{+}_{\alpha}{\mathfrak{V}}^{++}=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}{\mathfrak{V}}^{++}=0~{},$ (3.21)
is ill-defined since the exponential is invariant under the gauge transformations
${\mathfrak{V}}^{++}\to{\mathfrak{V}}^{++}+D^{++}\sigma~{},\qquad D^{+}_{\alpha}\sigma=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}\sigma=0~{}.$ (3.22)
This is the gauge freedom for the Abelian vector multiplet, and it can be fixed by applying the Faddeev-Popov procedure to result with
$\hat{\delta}\big{[}H^{++}\big{]}$ $=$ $\int\big{[}\mathscr{D}{\mathfrak{V}}^{++}\big{]}\,\delta\big{[}D^{++}{\mathfrak{V}}^{++}\big{]}\Delta_{\rm VM}\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}H^{++}\right\}~{},$ (3.23a)
$\Delta_{\rm VM}$ $=$ $\int\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,\tilde{C}(D^{++})^{2}C\right\}={\rm Det}\Big{[}(D^{++})^{2}\Big{]}~{}.$ (3.23b)
Here $\Delta_{\rm VM}$ denotes the Faddeev-Popov determinant arising in the vector multiplet model, and the Faddeev-Popov ghosts $\tilde{C}$ and $C$ are fermionic analytic superfields [66, 61] of the $\omega$-hypermultiplet type [60]. In accordance with [3], $\hat{\delta}\big{[}H^{++}\big{]}$ does not change if the gauge condition is deformed by $D^{++}{\mathfrak{V}}^{++}\to D^{++}{\mathfrak{V}}^{++}+\Omega^{(4)}$, where $\Omega^{(4)}$ is a background analytic superfield. Thus we obtain a more general representation for $\hat{\delta}\big{[}H^{++}\big{]}$:
$\hat{\delta}\big{[}H^{++}\big{]}$ $=$ $\int\big{[}\mathscr{D}{\mathfrak{V}}^{++}\big{]}\,\delta\big{[}D^{++}{\mathfrak{V}}^{++}-\Omega^{(4)}\big{]}\Delta_{\rm VM}\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}H^{++}\right\}~{}.$ (3.24)
Making use of the identity
$\delta\big{[}D^{++}{\mathfrak{V}}^{++}\big{]}=\int\big{[}\mathscr{D}\rho\big{]}\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}D^{++}\rho\right\}~{},$ (3.25)
another representation for $\hat{\delta}\big{[}H^{++}\big{]}$ can be obtained, which is:
$\hat{\delta}\big{[}H^{++}\big{]}$ $=$ $\int\big{[}\mathscr{D}\rho\big{]}{\delta}\big{[}H^{++}+D^{++}\rho\big{]}\Delta_{\rm VM}~{}.$ (3.26)
Relation (3.23) defines the generalised delta-function $\hat{\delta}\big{[}H^{++}\big{]}$ which should be used when trying to apply a modified version of the Faddeev-Poppov quantisation. Our next task is to compute a path integral over the Abelian gauge group (3.19)
$\int{\rm d}\mu_{{\mathfrak{U}}^{++}}\hat{\delta}\left[H^{++}-\frac{1}{8}(D^{+})^{2}\Lambda[{\mathfrak{U}}^{++}]-\frac{1}{8}(\bar{D}^{+})^{2}\bar{\Lambda}[{\mathfrak{U}}^{++}]\right]$ (3.27)
First of all, we need to identify a correct integration measure. Given a functional $F[\Psi,\bar{\Psi}]$, let us consider an integral
$I[\Psi,\bar{\Psi}]\propto\int\big{[}\mathscr{D}{\mathfrak{U}}^{++}\big{]}\,F\Big{[}\Psi+{\rm i}\Lambda[{\mathfrak{U}}^{++}],\bar{\Psi}-{\rm i}\bar{\Lambda}[{\mathfrak{U}}^{++}]\Big{]}\equiv\int\big{[}\mathscr{D}{\mathfrak{U}}^{++}\big{]}\,F\big{[}\Psi^{{\mathfrak{U}}^{++}},\bar{\Psi}^{{\mathfrak{U}}^{++}}\big{]}~{}.$ (3.28)
This integral is ill-defined since the integrand is invariant under the gauge transformations (3.22). Once again, this is the gauge freedom for the Abelian vector multiplet, and it can be fixed by applying the Faddeev-Popov procedure to result with
$I[\Psi,\bar{\Psi}]$ $=$ $\int{\rm d}\mu_{{\mathfrak{U}}^{++}}\,F\big{[}\Psi^{{\mathfrak{U}}^{++}},\bar{\Psi}^{{\mathfrak{U}}^{++}}\big{]}~{},$ (3.29)
${\rm d}\mu_{{\mathfrak{U}}^{++}}=\big{[}\mathscr{D}{\mathfrak{U}}^{++}\big{]}\delta\big{[}D^{++}{\mathfrak{U}}^{++}-\Upsilon^{(4)}\big{]}\Delta_{\rm VM}~{}.$
Here $\Upsilon^{(4)}$ is a background analytic superfield.
Now, were are prepared to evaluate
$(\Delta_{\rm TM})^{-1}$ ::=:= $\int{\rm d}\mu_{{\mathfrak{U}}^{++}}\,\hat{\delta}\left[H^{++}\big{(}\Psi^{{\mathfrak{U}}^{++}},\bar{\Psi}^{{\mathfrak{U}}^{++}}\big{)}\right]~{},$ (3.30)
where $H^{++}(\Psi,\bar{\Psi})$ denotes the gauge-fixing condition (3.12). Keeping in mind (3.24) and (3.29), this can be rewritten as
$(\Delta_{\rm TM})^{-1}$ $=$ $\int\big{[}\mathscr{D}{\mathfrak{V}}^{++}\big{]}\big{[}\mathscr{D}{\mathfrak{U}}^{++}\big{]}\delta\big{[}D^{++}{\mathfrak{V}}^{++}-\Omega^{(4)}\big{]}\delta\big{[}D^{++}{\mathfrak{U}}^{++}-\Upsilon^{(4)}\big{]}(\Delta_{\rm VM})^{2}$ (3.31)
×$\times\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}\Big{(}H^{++}-\frac{1}{8}(D^{+})^{2}\Lambda[{\mathfrak{U}}^{++}]-\frac{1}{8}(\bar{D}^{+})^{2}\bar{\Lambda}[{\mathfrak{U}}^{++}]\Big{)}\right\}~{}.$
Making use of the relation (3.20) allows us to rewrite the expression in second line as
$\exp\left\{{\rm i}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}\Big{(}H^{++}-\frac{1}{16}(D^{+})^{2}(\bar{D}^{+})^{2}{\mathfrak{U}}^{--}\Big{)}\right\}~{}.$ (3.32)
Next we perform the following shift in the path integral
${\mathfrak{U}}^{++}\to{\mathfrak{U}}^{++}+\frac{{\rm i}}{4}\Box^{-1}\Big{(}(D^{+})^{2}\Psi-(\bar{D}^{+})^{2}\bar{\Psi}\Big{)}~{},$ (3.33a)
which implies
${\mathfrak{U}}^{--}\to{\mathfrak{U}}^{--}+\frac{{\rm i}}{4}\Box^{-1}\Big{(}(D^{-})^{2}\Psi-(\bar{D}^{-})^{2}\bar{\Psi}\Big{)}~{}.$ (3.33b)
We also point out that
$\frac{1}{16}\int{\rm d}\zeta^{(-4)}\,{\mathfrak{V}}^{++}(D^{+})^{2}(\bar{D}^{+})^{2}{\mathfrak{U}}^{--}=\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{{\mathfrak{V}}^{++}(u_{1}){\mathfrak{U}}^{++}(u_{2})}{(u^{+}_{1}u^{+}_{2})^{2}}~{},$ (3.34)
where ${\rm d}^{4|8}z:={\rm d}^{4}x{\rm d}^{4}\theta{\rm d}^{4}\bar{\theta}$ is the integration measure of ${\cal N}=2$ Minkowski superspace. As a result, (3.31) turns into
$(\Delta_{\rm TM})^{-1}$ $=$ $\int\big{[}\mathscr{D}{\mathfrak{V}}^{++}\big{]}\big{[}\mathscr{D}{\mathfrak{U}}^{++}\big{]}\delta\big{[}D^{++}{\mathfrak{V}}^{++}-\Omega^{(4)}\big{]}\delta\big{[}D^{++}{\mathfrak{U}}^{++}-\Upsilon^{(4)}\big{]}(\Delta_{\rm VM})^{2}$ (3.35)
×$\times\exp\left\{-{\rm i}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{{\mathfrak{V}}^{++}(1){\mathfrak{U}}^{++}(2)}{(u^{+}_{1}u^{+}_{2})^{2}}\right\}~{}.$
Since the right-hand side is independent of the analytic superfields $\Omega^{(4)}$ and $\Upsilon^{(4)}$, we can integrate over them with a useful weight that we choose
$\Delta_{\rm NK}\exp\left\{{{\rm i}}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Omega^{(4)}(1)\,\frac{(u^{-}_{1}u^{-}_{2})}{(u^{+}_{1}u^{+}_{2})^{3}}\,\Upsilon^{(4)}(2)\right\}~{},$ (3.36)
with $\Delta_{\rm NK}$ the Nielsen-Kallosh determinant. The latter is determined by requiring
$1=\Delta_{\rm NK}\int\big{[}\mathscr{D}\Omega^{(4)}\big{]}\big{[}\mathscr{D}\Upsilon^{(4)}\big{]}\,\exp\left\{{{\rm i}}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Omega^{(4)}(1)\,\frac{(u^{-}_{1}u^{-}_{2})}{(u^{+}_{1}u^{+}_{2})^{3}}\,\Upsilon^{(4)}(2)\right\}~{}.$ (3.37)
One may evaluate $\Delta_{\rm NK}$ following the prescription given in [67] (see also [68] for a review) to end up with
$\Delta_{\rm NK}=\left(\Delta_{\rm VM}\right)^{-1}\,{\rm Det}_{(0,4)}\,\square~{}.$ (3.38)
Here we have introduced the functional determinant ${\rm Det}_{(0,4)}\,\square$ defined by
$\left({\rm Det}_{(0,4)}\,{\square}{}\right)^{-1}$ $=$ $\int[\mathscr{D}\rho^{(+4)}]\,[\mathscr{D}\sigma]\exp\left\{{\rm i}\,{\rm tr}\int{\rm d}\zeta^{(-4)}\,\rho^{(+4)}{\square}{}\,\sigma\right\}~{},$ (3.39a)
where $\rho^{(+4)}$ and $\sigma$ are unconstrained analytic superfields. In what follows, we will also need the following result
$\left({\rm Det}_{(2,2)}\,{\Box}{}\right)^{-1}$ $=$ $\int[\mathscr{D}U^{++}]\,[\mathscr{D}V^{++}]\,\exp\left\{{\rm i}\,{\rm tr}\int{\rm d}\zeta^{(-4)}\,U^{++}\square\,V^{++}\right\}~{},$ (3.39b)
The result of evaluating $\Delta_{\rm TM}$ is
$\Delta_{\rm TM}=(\Delta_{\rm NK})^{-1}\,(\Delta_{\rm VM})^{-2}\,{\rm Det}_{(2,2)}\,{\square}{}=(\Delta_{\rm VM})^{-1}\,\left({\rm Det}_{(0,4)}\,\square\right)^{-1}\,{\rm Det}_{(2,2)}\,\square{}~{}.$ (3.40)
At this stage, all prerequisites have been derived in order to quantise the tensor multiplet model (3.6). We start with the formal expression
$\int\big{[}\mathscr{D}\Psi\big{]}\big{[}\mathscr{D}\bar{\Psi}\big{]}\exp\left\{{\rm i}S[\Psi,\bar{\Psi}]\right\}$ (3.41)
and insert under the integral the following unit
11 11 $=$ $\Delta_{\rm TM}\int{\rm d}\mu_{{\mathfrak{U}}^{++}}\,\hat{\delta}\left[H^{++}\big{(}\Psi^{{\mathfrak{U}}^{++}},\bar{\Psi}^{{\mathfrak{U}}^{++}}\big{)}\right]$ (3.42)
$=$ $\Delta_{\rm TM}\int{\rm d}\mu_{{\mathfrak{U}}^{++}}\,\big{[}\mathscr{D}\rho\big{]}{\delta}\big{[}H^{++}+D^{++}\rho-\Xi^{++}\big{]}\Delta_{\rm VM}~{},$
where $\Xi^{++}$ is a background analytic superfield. By construction, (3.42) is independent of $\Xi^{++}$. After inserting (3.42) into (3.41), we make the change of variables $\Psi\to\Psi^{-{\mathfrak{U}}^{++}}$. As a result, the infinite group volume $\int{\rm d}\mu_{{\mathfrak{U}}^{++}}$ factorises and we end up with the partition function
$Z=\int\big{[}\mathscr{D}\Psi\big{]}\big{[}\mathscr{D}\bar{\Psi}\big{]}\big{[}\mathscr{D}\rho\big{]}{\delta}\big{[}H^{++}+D^{++}\rho-\Xi^{++}\big{]}\Delta_{\rm TM}\Delta_{\rm VM}\exp\left\{{\rm i}S[\Psi,\bar{\Psi}]\right\}~{}.$ (3.43)
It remains to integrate the right-hand side over $\Xi^{++}$ with a convenient weight
$\exp\left\{\frac{{\rm i}}{2}\int{\rm d}\zeta^{(-4)}\,(\Xi^{++})^{2}\right\}~{}.$
This leads to
$Z=\int\big{[}\mathscr{D}\Psi\big{]}\big{[}\mathscr{D}\bar{\Psi}\big{]}\big{[}\mathscr{D}\rho\big{]}\Delta_{\rm TM}\Delta_{\rm VM}\exp\left\{\frac{{\rm i}}{2}\int{\rm d}^{4|8}z\,\bar{\Psi}\Psi+\frac{{\rm i}}{2}\int{\rm d}\zeta^{(-4)}\,(D^{++}\rho)^{2}\right\}~{}.$ (3.44)
Evaluating the path integral over $\Psi$ and $\bar{\Psi}$ gives
$\int\big{[}\mathscr{D}\Psi\big{]}\big{[}\mathscr{D}\bar{\Psi}\big{]}\exp\left\{\frac{{\rm i}}{2}\int{\rm d}^{4|8}z\,\bar{\Psi}\Psi\right\}=\left({\rm Det}\,{\cal H}\right)^{-1/2}~{},$ (3.45)
where we have introduced the following operator
$\mathcal{H}=\begin{pmatrix}0&\bar{D}^{4}\\ D^{4}&0\end{pmatrix}~{},\qquad\bar{D}^{4}=\frac{1}{48}\bar{D}^{ij}\bar{D}_{ij}~{},$ (3.46)
which acts on the space of chiral/antichiral column vectors $(\Psi,\bar{\Psi})^{\rm T}$. As a result, for the partition function we obtain
$Z=\Big{\{}\left({\rm Det}\,{\cal H}\right)^{-1/2}\left({\rm Det}_{(2,2)}\,{\square}{}\right)\,\left({\rm Det}_{(0,4)}\,{\square}{}\right)^{-1}\Big{\}}\Big{(}{\rm Det}\big{[}(D^{++})^{2}\big{]}\Big{)}^{-1/2}~{}.$ (3.47)
The obtained expression simplifies drastically by taking into account the following identity:
$\left({\rm Det}\,{\cal H}\right)^{-1/2}\left({\rm Det}_{(2,2)}\,{\square}{}\right)\,\left({\rm Det}_{(0,4)}\,{\square}{}\right)^{-1}=1~{}.$ (3.48)
This identity may be derived by making use of a change of variables, $\{\Psi,\bar{\Psi},\omega,\sigma\}\to\{V^{++},U^{++}\}$, defined by
$V^{++}$ $=$ $D^{++}\omega+G^{++}~{},$ (3.49a)
$U^{++}$ $=$ $D^{++}\sigma+H^{++}~{},$ (3.49b)
and its inverse. Here $G^{++}$ and $H^{++}$ are defined as in (3.7a) and (3.12), respectively. In a supergravity background, (3.48) is expected to be replaced by a topological invariant containing the ${\cal N}=2$ Gauss-Bonnet term [69].
Our final result for the partition function of the tensor multiplet is
$Z=\Big{(}{\rm Det}\,\big{[}(D^{++})^{2}\big{]}\Big{)}^{-1/2}~{}.$ (3.50)
This is exactly the partition function of the $\omega$ hypermultiplet. Our conclusion will become non-trivial if a supergravity background is turned on.
The main results of this section may be extended to models for a self-interacting tensor multiplet described in [64, 61]. This amounts to replacing the free Lagrangian in (3.6) with an interacting one,
$\frac{1}{2}(G^{++})^{2}~{}\to~{}{\cal L}^{(+4)}\Big{(}G^{++},u^{\pm}\Big{)}~{}.$ (3.51)
4 Concluding comments
In section 2 we studied quantum equivalence of the super FT model (2.1) and the nonlinear $\sigma$-model (2.11). This equivalence may be immediately extended to the more general class of theories defined by replacing the $V^{2}$ term in (2.1) with a general function of $V$,
$S[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\text{tr}\int{\rm d}^{4|4}z\,\mathfrak{F}(V)~{},$ (4.1)
Specifically, by fixing $\mathfrak{F}(V)$ and repeating the same analysis as in section 2, one may show quantum equivalence to the following $\sigma$-model
$S[\Phi,\bar{\Phi}]=\text{tr}\int{\rm d}^{4|4}z\,\mathfrak{F}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (4.2)
Additionally, one may instead consider the model obtained by replacing the $V^{2}$ term in (2.1) with the super Yang-Mills action
$S_{\text{T}}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\Big{\{}\chi^{\alpha}W_{\alpha}+\frac{1}{g^{2}}W^{2}\Big{\}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{g^{2}}\bar{W}^{2}\Big{\}}~{}.$ (4.3)
In addition to invariance under (2.5), this model enjoys super Yang-Mills gauge symmetry [^10] We recall that $L_{A}B=[A,B]$, for operators $A$ and $B$.
$\delta_{\Lambda}V=-{{\rm i}}L_{V}(\Lambda+\bar{\Lambda})+{{\rm i}}L_{V}\coth(L_{V})(\bar{\Lambda}-\Lambda)~{},\qquad\bar{D}_{\dot{\alpha}}\Lambda=0~{}.$ (4.4)
Further, it is classically topological; it describes trivial dynamics. This property proves to also hold at the quantum level, which may be seen via a slight modification of the analysis of section 2. Specifically, it is necessary to incorporate the ghost action corresponding to (4.4) and choose a gauge fermion $\Psi(\mbox{\boldmath$\Phi$})$ such that the gauge $D^{2}V=\bar{D}^{2}V=0$ is enforced. With these adjustments, a similar analysis to that conducted in section 2 leads to $Z=1$ for this theory, implying trivial dynamics.
An interesting open problem is to extend the results of section 3 to a curved supergravity background. There are two general superspace approaches to formulate off-shell supergravity-matter systems in four dimensions. The harmonic superspace approach offers powerful prepotential formulations for ${\cal N}=2$ supergravity [70, 71] (reviewed in [61, 72]). The projective superspace approach proves to be ideal for developing covariant geometric formulations for supergravity-matter systems with eight supercharges [73, 74, 75, 76]. With the advent of ${\cal N}=2$ conformal superspace [77], and its applications to component reduction [78], a novel formulation of curved projective superspace has been given in [79, 80]. This approach has also been extended to a novel covariant harmonic superspace framework in [81]. For a review of covariant superspace approaches to ${\cal N}=2$ supergravity, see [82]. It appears that a combination of harmonic and projective superspace methods is required in order to extend the results of section 3 to ${\cal N}=2$ supergravity.
Covariant ${\cal N}=2$ supergravity techniques are expected to be indispensable for computing one-loop effective actions in curved superspace. For instance, the operator (3.46) will turn into
$\mathcal{H}=\begin{pmatrix}0&\bar{\Delta}\\ \Delta&0\end{pmatrix}~{},$ (4.5)
where $\bar{\Delta}$ denotes the chiral projecting operator [75, 83]
$\bar{\Delta}$ $=$ $\frac{1}{96}\Big{(}({\bar{\cal D}}^{ij}+16\bar{S}^{ij}){\bar{\cal D}}_{ij}-({\bar{\cal D}}^{{\dot{\alpha}}{\dot{\beta}}}-16\bar{Y}^{{\dot{\alpha}}{\dot{\beta}}}){\bar{\cal D}}_{{\dot{\alpha}}{\dot{\beta}}}\Big{)}$ (4.6)
$=$ $\frac{1}{96}\Big{(}{\bar{\cal D}}_{ij}({\bar{\cal D}}^{ij}+16\bar{S}^{ij})-{\bar{\cal D}}_{{\dot{\alpha}}{\dot{\beta}}}({\bar{\cal D}}^{{\dot{\alpha}}{\dot{\beta}}}-16\bar{Y}^{{\dot{\alpha}}{\dot{\beta}}})\Big{)}~{}.$
Here ${\cal D}_{A}=({\cal D}_{a},{\cal D}_{\alpha}^{i},\bar{\cal D}_{i}^{\dot{\alpha}})$ are the covariant derivatives in curved superspace, ${\bar{\cal D}}^{{\dot{\alpha}}{\dot{\beta}}}:={\bar{\cal D}}^{({\dot{\alpha}}}_{k}{\bar{\cal D}}^{{\dot{\beta}})k}$, and $\bar{S}^{ij}$ and $\bar{Y}^{{\dot{\alpha}}{\dot{\beta}}}$ are certain torsion tensors, see [82] for a review. The main properties of $\bar{\Delta}$ include the following: for any scalar $U$, it holds that
${\bar{\cal D}}^{{\dot{\alpha}}}_{i}\bar{\Delta}U$ $=$ $0~{},$ (4.7a)
$\int{\rm d}^{4}x\,{\rm d}^{4}\theta\,{\rm d}^{4}{\bar{\theta}}\,E\,U$ $=$ $\int{\rm d}^{4}x\,{\rm d}^{4}\theta\,{\cal E}\,\bar{\Delta}U~{},$ (4.7b)
where $E$ and ${\cal E}$ denote, respectively, the full superspace and the chiral subspace densities. Operator $\bar{\Delta}\Delta$ is a covariantly chiral fourth-order operator when acting on the space of covariantly chiral scalar superfields,
$\bar{\cal D}^{\dot{\alpha}}_{i}\Psi=0\quad\implies\quad\bar{\Delta}\Delta\Psi=\Big{\{}({\cal D}^{a}{\cal D}_{a})^{2}+\dots\Big{\}}\Psi~{}.$ (4.8)
In the supersymmetric literature, the BV procedure has been employed [84, 85, 86, 87] to quantise theories involving the massless non-minimal scalar multiplet [88] described by a complex linear superfield. These are reducible gauge theories of infinite stage of reducibility, which is similar to the Green-Schwarz superstring. Strictly speaking, the BV approach is directly applicable to quantise finitely reducible gauge theories. However, a way out was found in Refs. [84, 85] and their extensions [86, 87]. Another family of reducible supersymmetric theories are the massless higher-spin supermultiplets in AdS4 [89, 90]. When realised in terms of unconstrained superfield prepotentials introduced in [89, 91, 92, 90], their stage of reducibility is finite or infinite depending on the superspin and the type of the formulation. Lagrangian quantisation of the higher-spin supermultiplets in AdS4 [89] was carried out in [93] using a special simplification of the BV procedure in reduction coordinates (transversal irreducible superfields) for a general quadratic action. It would be interesting to come back to the problem of BV quantisation of the supersymmetric higher-spin gauge models in AdS4 in terms of unconstrained superfields.
Acknowledgements: This work was supported in part by the Australian Research Council, project No. DP230101629. Since May 1, 2024, the work of ER is supported by the Brian Dunlop Physics Fellowship. We acknowledge the kind hospitality and financial support extended to us at the MATRIX Program “New Deformations of Quantum Field and Gravity Theories” (between 22 January and 2 February, 2024) where part of this work was performed.
Appendix A Rudiments of harmonic superspace
The ${\cal N}=2$ harmonic superspace ${\mathbb{R}}^{4|8}\times S^{2}$ [60, 61] extends conventional ${\cal N}=2$ superspace ${\mathbb{R}}^{4|8}$ (paramerised by coordinates $z^{A}=(x^{a},\theta^{\alpha}_{i},{\bar{\theta}}_{\dot{\alpha}}^{i})$, with $i=\hat{1},\hat{2}$) by the two-sphere $S^{2}=\mathsf{SU}(2)/\mathsf{U}(1)$ parametrised by harmonics elements
$({u_{i}}^{-}\,,\,{u_{i}}^{+})\in\mathsf{SU}(2)~{},\quad u^{+}_{i}=\varepsilon_{ij}u^{+j}~{},\quad\overline{u^{+i}}=u^{-}_{i}~{},\quad u^{+i}u_{i}^{-}=1~{}.$ (A.1)
Instead of using the standard basis for spinor covariant derivatives $D^{i}_{\alpha}$ and ${\bar{D}}^{\dot{\alpha}}_{i}$, with $i=\hat{1},\hat{2}$, which obey the anti-commutation relations
$\{D^{i}_{\alpha},D^{j}_{\beta}\}=\{{\bar{D}}^{\dot{\alpha}}_{i},{\bar{D}}^{\dot{\beta}}_{j}\}=0~{},\qquad\{D^{i}_{\alpha},{\bar{D}}_{{\dot{\beta}}j}\}=-2{\rm i}\,\delta^{i}_{j}\,(\sigma^{c})_{\alpha{\dot{\beta}}}\,\partial_{c}~{},$ (A.2)
a new harmonic-dependent basis can be introduced by the rule
$D^{\pm}_{\alpha}=D^{i}_{\alpha}\,u^{\pm}_{i}~{},\qquad{\bar{D}}^{\pm}_{\dot{\alpha}}={\bar{D}}^{i}_{\dot{\alpha}}\,u^{\pm}_{i}~{}.$ (A.3)
The operators $D^{+}_{\alpha}$ and $\bar{D}^{+}_{\dot{\alpha}}$ strictly anticommute, and therefore on can introduce analytic superfields of $\mathsf{U}(1)$ charge $n$, $\varphi^{(n)}(z,u)$, with the properties:
$D^{+}_{\alpha}\varphi^{(n)}=0~{},\qquad\bar{D}^{+}_{\dot{\alpha}}\varphi^{(n)}=0;$ (A.4a)
$\varphi^{(n)}(z,{\rm e}^{{\rm i}\alpha}\,u^{+},{\rm e}^{-{\rm i}\alpha}\,u^{-})={\rm e}^{{\rm i}n\alpha}\,\varphi^{(n)}(z,u^{+},u^{-})~{}.$ (A.4b)
Such superfields live on the so-called analytic subspace of the harmonic superspace. When working in the harmonic-superspace approach, of special significance are the operators
$D^{++}=u^{+i}\frac{\partial}{\partial u^{-i}}~{},\qquad D^{--}=u^{-i}\frac{\partial}{\partial u^{+i}}~{}.$ (A.5)
Using the fourth-order operators
$(D^{+})^{4}={1\over 16}(D^{+})^{2}({\bar{D}}^{+})^{2}~{},\qquad(D^{-})^{4}={1\over 16}(D^{-})^{2}({\bar{D}}^{-})^{2}~{},$ (A.6)
integration over the analytic subspace is defined by
$\int{\rm d}\zeta^{(-4)}\,L^{(+4)}=\int{\rm d}^{4}x\int{\rm d}u\,(D^{-})^{4}L^{(+4)}~{},\qquad D^{+}_{\alpha}L^{(+4)}={\bar{D}}^{+}_{\dot{\alpha}}L^{(+4)}=0~{}.$ (A.7)
Additionally, integration over the group manifold $\mathsf{SU}(2)$ is defined according to [60]
[…]

### v2

@@
 Email: sergei.kuzenko@uwa.edu.au, emmanouil.raptakis@uwa.edu.au
-The Batalin-Vilkovisky formalism is applied to quantise the ${\cal N}=1$ supersymmetric generalisation of the Freedman-Townsend (FT) model proposed by Lindström and Roček in 1983.
+The Batalin-Vilkovisky formalism is applied to quantise the ${\cal N}=1$ supersymmetric generalisation of the Freedman-Townsend (FT) model, which was proposed by Lindström and Roček in 1983 in Minkowski superspace and is lifted to a supergravity background in this paper.
 This super FT theory describes a non-Abelian tensor multiplet and is known to be classically equivalent to a supersymmetric nonlinear sigma model.
-Using path integral considerations, we demonstrate that this equivalence holds at the quantum level.
+Using path integral considerations, we demonstrate that this equivalence holds at the quantum level in the sense that the quantum supercurrents in the two theories coincide.
 A modified Faddeev-Popov procedure is employed to quantise models for the ${\cal N}=2$ tensor multiplet in harmonic superspace.
+The obtained results agree with those derived by applying the Batalin-Vilkovisky scheme within the harmonic superspace setting.
 1 Introduction
@@
 This is a non-Abelian extension of the ${\cal N}=1$ tensor multiplet proposed by Siegel [41].
+In this paper we will couple the ${\cal N}=1$ super FT model to supergravity.
 Covariant quantisation of the free tensor multiplet model in a supergravity background was carried out in [23, 42].
@@
 This paper is organised as follows.
-Section 2 is devoted to: (i) carrying out BV quantisation of the ${\cal N}=1$ super FT model (2.1); and (ii) proving its quantum equivalence to the supersymmetric nonlinear $\sigma$-model (2.11).
+Section 2 is devoted to: (i) coupling the super FT model to old minimal supergravity; (ii) carrying out BV quantisation of the ${\cal N}=1$ super FT model (2.1); and (iii) proving its quantum equivalence to the locally supersymmetric nonlinear $\sigma$-model (2.13).
 Section 3 is devoted to covariant quantisation of ${\cal N}=2$ tensor multiplet models.
@@
 Rudiments of harmonic superspace are collected in appendix A.
+BV quantisation of the ${\cal N}=2$ Abelian vector and tensor multiplets is carried out in appendix B.
 2 Supersymmetric Freedman-Townsend model
@@
 We also make use of the notational shorthands ${\rm d}^{4|4}z:={\rm d}^{4}x{\rm d}^{2}\theta{\rm d}^{2}\bar{\theta}$, ${\rm d}^{4|2}z_{+}:={\rm d}^{4}x{\rm d}^{2}\theta$ and ${\rm d}^{4|2}z_{-}:={\rm d}^{4}x{\rm d}^{2}\bar{\theta}$ to denote the integration measures of the Minkowski superspace and its chiral and antichiral subspaces.
-$S_{0}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,V^{2}~{},$ (2.1)
+$S_{0}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,E\,V^{2}~{},$ (2.1)
 where $\chi_{\alpha}$, $W_{\alpha}$ and $V={V}^{\dagger}$ are Lie algebra-valued superfields [^6] We will say that $\Upsilon$ is a Lie algebra-valued superfield if $\Upsilon=\Upsilon^{I}T^{I}$, where $\Upsilon^{I}$ is a superfield and $T^{I}$ are the generators of a semi-simple Lie group $G$.
@@
 satisfying:
-¯$\bar{\nabla}_{\dot{\alpha}}\chi_{\alpha}=0~{},\qquad\bar{\nabla}_{{\dot{\alpha}}}W_{\alpha}=0~{},\qquad\nabla^{\alpha}W_{\alpha}=\bar{\nabla}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}~{}.$ (2.2)
-Operators $\nabla_{A}=(\nabla_{a},\nabla_{\alpha},\bar{\nabla}^{\dot{\alpha}})$ above are gauge-covariant derivatives and obey the algebra:
-$\{\nabla_{\alpha},\nabla_{\beta}\}=0~{},\quad$$\{\nabla_{\alpha},\bar{\nabla}_{\dot{\alpha}}\}=-2{\rm i}\nabla_{\alpha{\dot{\alpha}}}~{},\quad\{\bar{\nabla}_{\dot{\alpha}},\bar{\nabla}_{\dot{\beta}}\}=0~{},$ (2.3a)
-$[\nabla_{\alpha},\nabla_{\beta{\dot{\beta}}}]$ $={\rm i}\varepsilon_{\alpha\beta}\bar{W}_{\dot{\alpha}}~{},\qquad[\bar{\nabla}_{\dot{\alpha}},\nabla_{\beta{\dot{\beta}}}]={\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}W_{\alpha}~{},$ (2.3b)
-$[\nabla_{\alpha{\dot{\alpha}}},\nabla_{\beta{\dot{\beta}}}]$ $=-\frac{1}{2}\varepsilon_{\alpha\beta}\bar{\nabla}_{({\dot{\alpha}}}\bar{W}_{{\dot{\beta}})}-\frac{1}{2}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\nabla_{(\alpha}W_{\beta)}~{}.$ (2.3c)
-These constraints may be solved [45] in terms of the complex unconstrained prepotential $\Omega=\Omega^{I}T^{I}$ and its Hermitian conjugate $\bar{\Omega}\equiv\Omega^{\dagger}$,
-$\nabla_{\alpha}={\rm e}^{-\Omega}D_{\alpha}{\rm e}^{\Omega}~{},\qquad\bar{\nabla}_{\dot{\alpha}}={\rm e}^{\bar{\Omega}}\bar{D}_{\dot{\alpha}}{\rm e}^{-\bar{\Omega}}~{},\qquad{\rm e}^{V}={\rm e}^{\Omega}{\rm e}^{\bar{\Omega}}~{},$ (2.4)
-where $D_{A}=(\partial_{a},D_{\alpha},\bar{D}^{\dot{\alpha}})$ are the covariant derivatives of Minkowski superspace.
-[^7] Our normalisation of the gauge prepotentials in (2.4) differs from [44] and agrees with [46].
+¯$\bar{\nabla}_{\dot{\alpha}}\chi_{\alpha}=0~{},\qquad\bar{\nabla}_{{\dot{\alpha}}}W_{\alpha}=0~{},\qquad\nabla^{\alpha}W_{\alpha}=\bar{\nabla}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}~{},$ (2.2)
+and $E^{-1}={\rm Ber}(E_{A}{}^{M})$ is the full superspace measure, while $\mathcal{E}$ denotes the chiral density.
+Operators $\nabla_{A}=(\nabla_{a},\nabla_{\alpha},\bar{\nabla}^{\dot{\alpha}})$ above are gauge-covariant derivatives, $\nabla_{A}=E_{A}{}^{M}\partial_{M}+\frac{1}{2}\Omega_{A}{}^{bc}M_{bc}+{\rm i}\,\Gamma_{A}{}^{I}T^{I}$, with $M_{bc}$ the Lorentz generators.
+The covariant derivatives obey the algebra:
+$\{\nabla_{\alpha},\nabla_{\beta}\}$ $=-4\bar{R}M_{\alpha\beta}~{},\quad\{\nabla_{\alpha},\bar{\nabla}_{\dot{\alpha}}\}=-2{\rm i}\nabla_{\alpha{\dot{\alpha}}}~{},\quad\{\bar{\nabla}_{\dot{\alpha}},\bar{\nabla}_{\dot{\beta}}\}=4R\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{},$ (2.3a)
+[¯$[\bar{\nabla}_{\dot{\alpha}},\nabla_{\beta{\dot{\beta}}}]$ $=-{\rm i}\varepsilon_{{\dot{\alpha}}{\dot{\beta}}}\Big{(}R\nabla_{\beta}+G_{\beta}{}^{{\dot{\gamma}}}\bar{\nabla}_{\dot{\gamma}}-\bar{\nabla}^{\dot{\gamma}}G_{\beta}{}^{{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+2W_{\beta}{}^{\gamma\delta}M_{\gamma\delta}-{W}_{\beta}\Big{)}-{\rm i}\nabla_{\beta}R\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{},$ (2.3b)
+$[{\nabla}_{\alpha},\nabla_{\beta{\dot{\beta}}}]$ $={\rm i}\varepsilon_{\alpha\beta}\Big{(}\bar{R}\bar{\nabla}_{\dot{\beta}}+G^{\gamma}{}_{{\dot{\beta}}}{\nabla}_{\gamma}-{\nabla}^{\gamma}G^{\delta}{}_{\dot{\beta}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+2\bar{W}_{\dot{\beta}}{}^{{\dot{\gamma}}{\dot{\delta}}}\bar{M}_{{\dot{\gamma}}{\dot{\delta}}}+\bar{W}_{\dot{\beta}}\Big{)}+{\rm i}\bar{\nabla}_{{\dot{\beta}}}\bar{R}\bar{M}_{{\dot{\alpha}}{\dot{\beta}}}~{}.$ (2.3c)
+Here the torsion tensors $R$, $G_{a}={\bar{G}}_{a}$ and $W_{\alpha\beta\gamma}=W_{(\alpha\beta\gamma)}$ satisfy the Bianchi identities:
+¯$\bar{\nabla}_{\dot{\alpha}}R=0~{},~{}~{}~{}~{}~{}~{}\bar{\nabla}_{\dot{\alpha}}W_{\alpha\beta\gamma}=0~{},$ (2.4a)
+¯$\bar{\nabla}^{\dot{\gamma}}G_{\alpha{\dot{\gamma}}}=\nabla_{\alpha}R~{},~{}~{}~{}~{}~{}~{}\nabla^{\gamma}W_{\alpha\beta\gamma}={\rm i}\,\nabla_{(\alpha}{}^{\dot{\gamma}}G_{\beta){\dot{\gamma}}}~{}.$ (2.4b)
+In this supergeometry integrals over the full superspace and its chiral subspace are related via:
+$\int{\rm d}^{4|4}z\,E\,U=-\frac{1}{4}\int{\rm d}^{4|2}z_{+}\,\,{\cal E}\,(\bar{\nabla}^{2}-4R)U~{}.$ (2.5)
+Constraints (2.3) may be solved [45] in terms of the complex unconstrained prepotential $\Omega=\Omega^{I}T^{I}$ and its Hermitian conjugate $\bar{\Omega}\equiv\Omega^{\dagger}$,
+$\nabla_{\alpha}={\rm e}^{-\Omega}\mathfrak{D}_{\alpha}{\rm e}^{\Omega}~{},\qquad\bar{\nabla}_{\dot{\alpha}}={\rm e}^{\bar{\Omega}}\bar{\mathfrak{D}}_{\dot{\alpha}}{\rm e}^{-\bar{\Omega}}~{},\qquad{\rm e}^{V}={\rm e}^{\Omega}{\rm e}^{\bar{\Omega}}~{},$ (2.6)
+where $\mathfrak{D}_{A}=(\mathfrak{D}_{a},\mathfrak{D}_{\alpha},\bar{\mathfrak{D}}^{\dot{\alpha}})$ are the covariant derivatives of the GWZ superspace [46], see e.g.
+[44, 47] for a review.
+[^7] Relation of the GWZ superspace geometry to other ${\cal N}=1$ supergravity formulations in spelled out in [48].
+Our normalisation of the gauge prepotentials in (2.6) differs from [44] and agrees with [47].
 The superfields $\bar{\chi}_{\dot{\alpha}}$ and $\bar{W}_{\dot{\alpha}}$ in (2.1) are Hermitian conjugates of $\chi_{\alpha}$ and $W_{\alpha}$, respectively, in the vector representation.
@@
 Making use of (2.2), it follows that action (2.1) is invariant under the following gauge transformation
-$\delta_{L}\chi_{\alpha}=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}L~{},\qquad L=\bar{L}~{},$ (2.5)
+$\delta_{L}\chi_{\alpha}=-\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}L~{},\qquad L=\bar{L}~{},$ (2.7)
 where the gauge parameter $L=L^{I}T^{I}$ is Hermitian but otherwise unconstrained.
@@
 The tensor multiplet is known to admit self-interactions of the form [41, 37]
-$S=\int{\rm d}^{4|4}z\,{\mathfrak{F}}(G)~{},\qquad G=\frac{1}{2}(D^{\alpha}\eta_{\alpha}+\bar{D}_{\dot{\alpha}}\bar{\eta}^{\dot{\alpha}})~{},\qquad\bar{D}_{\dot{\beta}}\eta_{\alpha}=0~{}.$ (2.7)
+$S=\int{\rm d}^{4|4}z\,E\,{\mathfrak{F}}(G)~{},\qquad G=\frac{1}{2}(\mathfrak{D}^{\alpha}\eta_{\alpha}+\bar{\mathfrak{D}}_{\dot{\alpha}}\bar{\eta}^{\dot{\alpha}})~{},\qquad\bar{\mathfrak{D}}_{\dot{\beta}}\eta_{\alpha}=0~{}.$ (2.9)
 Here $\eta_{\alpha}$ denotes the chiral prepotential of a single tensor multiplet, to distinguish this case from the super FT model.
@@
 The model has a classically equivalent (or dual) formulation [41, 37] realised in terms of a chiral scalar $\varphi$ and its conjugate $\bar{\varphi}$ with the action
-$S=\int{\rm d}^{4|4}z\,{\mathfrak{U}}(\varphi+\bar{\varphi})~{},\qquad\bar{D}_{\dot{\beta}}\varphi=0~{},$ (2.8)
+$S=\int{\rm d}^{4|4}z\,E\,{\mathfrak{U}}(\varphi+\bar{\varphi})~{},\qquad\bar{\mathfrak{D}}_{\dot{\beta}}\varphi=0~{},$ (2.10)
 where ${\mathfrak{U}}$ is the Legendre transform of ${\mathfrak{F}}$ [37],
@@
 To see this, we first note that the dynamical equations for $\chi_{\alpha}$ and its conjugate imply that $V$ is pure gauge
-$\frac{\delta_{l}S_{0}[\chi,\bar{\chi},V]}{\delta\chi^{\alpha}}=\frac{1}{2}W_{\alpha}=0\quad\implies\quad V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.10)
+$\frac{\delta_{l}S_{0}[\chi,\bar{\chi},V]}{\delta\chi^{\alpha}}=\frac{1}{2}W_{\alpha}=0\quad\implies\quad V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Phi=0~{}.$ (2.12)
 As a result, the action (2.1) reduces to
-$S[\Phi,\bar{\Phi}]=\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,\log^{2}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (2.11)
+$S[\Phi,\bar{\Phi}]=\frac{1}{2}\text{tr}\int{\rm d}^{4|4}z\,E\,\log^{2}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (2.13)
 This is a supersymmetric nonlinear $\sigma$-model with Kähler potential
@@
 One of the goals of this section, which we pursue below, is to prove that this equivalence extends to the quantum level.
-Specifically, our goal is to demonstrate that the partition functions for models (2.1) and (2.11) coincide.
+Specifically, our goal is to demonstrate that the partition functions for models (2.1) and (2.13) differ by a topological invariant.
+This means that the quantum supercurrents in these theories coincide.
 2.2 Quantisation via Batalin-Vilkovisky formalism
@@
 (2.7):
-$\delta_{L}\chi^{\alpha I}(z)$ $=\int{\rm d}^{4|4}z\,{R}^{\alpha I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}^{\alpha I\,J}(z,z^{\prime})=-\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla^{\alpha}\delta^{8}(z-z^{\prime})\delta^{IJ}~{},$ (2.14a)
-$\delta_{L}\bar{\chi}_{\dot{\alpha}}^{I}(z)$ $=\int{\rm d}^{4|4}z\,{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})=\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\delta^{8}(z-z^{\prime})\delta^{IJ}~{}.$ (2.14b)
+$\delta_{L}\chi^{\alpha I}(z)$ $=\int{\rm d}^{4|4}z\,E\,{R}^{\alpha I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}^{\alpha I\,J}(z,z^{\prime})=-\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\delta^{8}(z,z^{\prime})\delta^{IJ}~{},$ (2.16a)
+$\delta_{L}\bar{\chi}_{\dot{\alpha}}^{I}(z)$ $=\int{\rm d}^{4|4}z\,E\,{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})L^{J}(z^{\prime})\quad\implies\quad{R}_{\dot{\alpha}}^{I\,J}(z,z^{\prime})=\frac{{\rm i}}{4}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\delta^{8}(z,z^{\prime})\delta^{IJ}~{}.$ (2.16b)
 We also recall that, in the present model, the gauge parameters $L^{I}$ possess their own gauge symmetry, see eq.
@@
 In the present case, the nontrivial null eigenvectors $Z^{\underline{\alpha}}{}_{\underline{\alpha}_{1}}$ of the generators are:
-$\delta_{\sigma}L^{I}(z)$ $=\int{\rm d}^{4|2}z_{+}\,Z^{I\,J}(z,z^{\prime})\sigma(z^{\prime})+\int{\rm d}^{4|2}z_{-}\,\bar{Z}^{I\,J}(z,z^{\prime})\bar{\sigma}(z^{\prime})~{},$ (2.16a)
+$\delta_{\sigma}L^{I}(z)$ $=\int{\rm d}^{4|2}z_{+}\,{\cal E}\,Z^{I\,J}(z,z^{\prime})\sigma^{J}(z^{\prime})+\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\bar{Z}^{I\,J}(z,z^{\prime})\bar{\sigma}^{J}(z^{\prime})~{},$ (2.18a)
 $\implies$ $\quad Z^{I\,J}(z,z^{\prime})=\delta_{+}(z,z^{\prime})\delta^{IJ}~{},\qquad\bar{Z}^{I\,J}(z,z^{\prime})=\delta_{-}(z,z^{\prime})\delta^{IJ}~{}.$ (2.18b)
@@
 Then, the in-out vacuum amplitude for $S_{0}[\varphi]$ may be defined as
-$Z=\int\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{\underline{A}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}\big{]}\,\mu\,\delta\Big{[}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}-\frac{\delta\Psi}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\Big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]\Big{)}~{},$ (2.19)
+$Z=\int\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{\underline{A}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}\big{]}\,\mu\,\delta\Big{[}\mbox{\boldmath$\Phi$}^{*}_{\underline{A}}-\frac{\delta\Psi}{\delta{\mbox{\boldmath$\Phi$}}^{\underline{A}}}\Big{]}\exp\Big{(}{\rm i}S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]\Big{)}~{},$ (2.21)
 where $\mu$ is the quantum integration measure and the fermionic functional $\Psi[\mbox{\boldmath$\Phi$}]$ is the gauge fermion, whose only requirement is that $Z$ possesses no residual gauge freedom.
@@
 Making use of equations (2.16), (2.18) and (2.19), we find $S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ to be:
-$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ $=S_{0}[\chi,\bar{\chi},V]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}{\rm i}\big{(}\nabla^{\alpha}\chi_{\alpha}^{*}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{*{\dot{\alpha}}}\big{)}C+C^{*}(\eta+\bar{\eta})\Big{\}}$
-$\phantom{=}+{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\chi^{*\alpha},\chi_{\alpha}^{*}\big{]}\eta-{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\bar{\chi}^{*}_{\dot{\alpha}},\bar{\chi}^{{\dot{\alpha}}*}\big{]}\bar{\eta}~{},$ (2.22)
+$S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]$ $=S_{0}[\chi,\bar{\chi},V]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}{\rm i}\big{(}\nabla^{\alpha}\chi_{\alpha}^{*}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{*{\dot{\alpha}}}\big{)}C+C^{*}(\eta+\bar{\eta})\Big{\}}$
+$\phantom{=}+{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}\chi^{*\alpha},\chi_{\alpha}^{*}\big{]}\eta-{\rm i}\,\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}\bar{\chi}^{*}_{\dot{\alpha}},\bar{\chi}^{{\dot{\alpha}}*}\big{]}\bar{\eta}~{},$ (2.24)
 where the ghost superfield $\eta$ is covariantly chiral, ¯$\bar{\nabla}_{\dot{\alpha}}\eta=0$.
@@
 These will be used to obtain a new solution of master equation (2.20) that is more useful for the elimination of antifields. It is then easily seen that the following functional
-$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]$ $=S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\tilde{C}^{*}\pi+\mbox{\boldmath$\eta$}^{*}\bar{\mbox{\boldmath$\pi$}}+\bar{\mbox{\boldmath$\eta$}}^{*}\mbox{\boldmath$\pi$}+\eta^{\prime*}\bar{\pi}^{\prime}+\bar{\eta}^{\prime*}\pi^{\prime}$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\Lambda^{*}\bar{\Pi}+\bar{\Lambda}^{*}\Pi+\Lambda^{\prime*}\bar{\Pi}^{\prime}+\bar{\Lambda}^{*}\Pi^{\prime}\Big{\}}$ (2.24)
+$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]$ $=S[\mbox{\boldmath$\Phi$}_{\text{min}},{\mbox{\boldmath$\Phi$}}^{*}_{\text{min}}]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\tilde{C}^{*}\pi+\frac{1}{2}\Big{[}(\mbox{\boldmath$\eta$}^{*}+\bar{\mbox{\boldmath$\pi$}})^{2}+(\bar{\mbox{\boldmath$\eta$}}^{*}+\mbox{\boldmath$\pi$})^{2}+(\eta^{\prime*}+\bar{\pi}^{\prime})^{2}$
+$\qquad+(\bar{\eta}^{\prime*}+\pi^{\prime})^{2}+(\Lambda^{*}+\bar{\Pi})^{2}+(\bar{\Lambda}^{*}+\Pi)^{2}+(\Lambda^{\prime*}+\bar{\Pi}^{\prime})^{2}+(\bar{\Lambda}^{*}+\Pi^{\prime})^{2}\Big{]}\Big{\}}$ (2.26)
 is also a solution to the master equation (2.20). It should be noted that this functional satisfies
@@
 therefore the quantum measure in (2.21) may be set to unity, $\mu=1$, in accordance with [49]. Choosing the gauge fermion to be
-$\Psi(\mbox{\boldmath$\Phi$})$ $=\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{{\rm i}}{2}\tilde{C}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+C(\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}})+\tilde{C}(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\tilde{C}\pi\Big{\}}$
-$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\Big{\{}-\frac{1}{n}\big{(}{\mbox{\boldmath$\eta$}}\pi^{\prime}+\mbox{\boldmath$\pi$}\eta^{\prime}\big{)}+\Lambda\Pi+\Lambda^{\prime}\Pi^{\prime}\Big{\}}$
-$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}-\frac{1}{n}\big{(}\bar{\mbox{\boldmath$\eta$}}\bar{\pi}^{\prime}+\bar{\mbox{\boldmath$\pi$}}\bar{\eta}^{\prime}\big{)}+\bar{\Lambda}\bar{\Pi}+\bar{\Lambda}^{\prime}\bar{\Pi}^{\prime}\Big{\}}~{},$ (2.26)
+$\Psi(\mbox{\boldmath$\Phi$})$ $=\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{{\rm i}}{2}\tilde{C}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+C(\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}})+\tilde{C}(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\tilde{C}\pi\Big{\}}$
+$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\Big{\{}-\frac{1}{n}\big{(}{\mbox{\boldmath$\eta$}}\pi^{\prime}+\mbox{\boldmath$\pi$}\eta^{\prime}\big{)}+\Lambda\Pi+\Lambda^{\prime}\Pi^{\prime}\Big{\}}$
+$\phantom{=}\quad+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\Big{\{}-\frac{1}{n}\big{(}\bar{\mbox{\boldmath$\eta$}}\bar{\pi}^{\prime}+\bar{\mbox{\boldmath$\pi$}}\bar{\eta}^{\prime}\big{)}+\bar{\Lambda}\bar{\Pi}+\bar{\Lambda}^{\prime}\bar{\Pi}^{\prime}\Big{\}}~{},$ (2.28)
 and eliminating antifields in eq. (2.2) leads to
-$S[\mbox{\boldmath$\Phi$},{\mbox{\boldmath$\Phi$}}^{*}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta-\bar{\Pi}\Pi-\bar{\Pi}^{\prime}\Pi^{\prime}$
-$\qquad\qquad\qquad+\Big{(}\frac{{\rm i}}{2}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\pi\Big{)}\pi$
-$\qquad\qquad\qquad+\Big{(}-\frac{1}{2n}\pi^{\prime}-\frac{1}{4}\bar{\nabla}^{2}C\Big{)}\bar{\mbox{\boldmath$\pi$}}+\Big{(}-\frac{1}{2n}\bar{\pi}^{\prime}-\frac{1}{4}\nabla^{2}C\Big{)}\mbox{\boldmath$\pi$}$
-$\qquad\qquad\qquad+\Big{(}\frac{1}{2n}\mbox{\boldmath$\pi$}-\frac{1}{4}\bar{\nabla}^{2}\tilde{C}\Big{)}\bar{\pi}^{\prime}+\Big{(}\frac{1}{2n}\bar{\mbox{\boldmath$\pi$}}-\frac{1}{4}\nabla^{2}\tilde{C}\Big{)}\pi^{\prime}\Big{\}}$
-$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.27)
+$S[\mbox{\boldmath$\Phi$},{\mbox{\boldmath$\Phi$}}^{*}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{1}{8}\tilde{C}\big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\big{)}C+\frac{1}{2}(\bar{\eta}+\mbox{\boldmath$\eta$})^{2}$
+$\phantom{=}\quad+\frac{1}{2}(\bar{\mbox{\boldmath$\eta$}}+\eta)^{2}-\frac{1}{2}(\bar{\Pi}+\Pi)^{2}-\frac{1}{2}(\bar{\Pi}^{\prime}+\Pi^{\prime})^{2}+\Big{(}\frac{{\rm i}}{2}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}+(\eta^{\prime}+\bar{\eta}^{\prime})-\frac{1}{2m}\pi\Big{)}\pi$
+$\phantom{=}\quad+\frac{1}{2}\Big{(}-\frac{1}{2n}\pi^{\prime}-\frac{1}{4}(\bar{\nabla}^{2}-4R)C+\bar{\mbox{\boldmath$\pi$}}\Big{)}^{2}+\frac{1}{2}\Big{(}-\frac{1}{2n}\bar{\pi}^{\prime}-\frac{1}{4}(\nabla^{2}-4\bar{R})C+\mbox{\boldmath$\pi$}\Big{)}^{2}$
+$\phantom{=}\quad+\frac{1}{2}\Big{(}\frac{1}{2n}\mbox{\boldmath$\pi$}-\frac{1}{4}(\bar{\nabla}^{2}-4R)\tilde{C}+\bar{\pi}^{\prime}\Big{)}^{2}+\frac{1}{2}\Big{(}\frac{1}{2n}\bar{\mbox{\boldmath$\pi$}}-\frac{1}{4}(\nabla^{2}-4\bar{R})\tilde{C}+\pi^{\prime}\Big{)}^{2}\Big{\}}$
+$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\tilde{C},(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\tilde{C}\big{]}\eta$
+$\phantom{=}\quad+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\tilde{C},(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.29)
 […]
-$Z$ $=\int\big{[}\mathscr{D}\mbox{\boldmath$\varphi$}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\varphi$}]\Big{)}~{},\qquad\mbox{\boldmath$\varphi$}=\Big{\{}\chi^{\alpha},\bar{\chi}_{\dot{\alpha}},V,C,\tilde{C},\eta,\bar{\eta},\mbox{\boldmath$\eta$},\bar{\mbox{\boldmath$\eta$}},\eta^{\prime},\bar{\eta}^{\prime}\Big{\}}~{},$
-$S[\mbox{\boldmath$\varphi$}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}+\frac{n}{2}\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta$
-$\qquad\qquad\qquad\qquad+m\bar{\eta}^{\prime}\eta^{\prime}-\frac{m}{8}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}^{2}\Big{\}}$
-$\phantom{=}~{}-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.28)
+$Z$ $=\int\big{[}\mathscr{D}\mbox{\boldmath$\varphi$}\big{]}\exp\Big{(}{\rm i}S[\mbox{\boldmath$\varphi$}]\Big{)}~{},\qquad\mbox{\boldmath$\varphi$}=\Big{\{}\chi^{\alpha},\bar{\chi}_{\dot{\alpha}},V,C,\tilde{C},\eta,\bar{\eta},\mbox{\boldmath$\eta$},\bar{\mbox{\boldmath$\eta$}},\eta^{\prime},\bar{\eta}^{\prime}\Big{\}}~{},$
+$S[\mbox{\boldmath$\varphi$}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{1}{8}\tilde{C}\Big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}$
+$\phantom{=}\quad+\frac{n}{2}\big{[}({\nabla}^{2}-4\bar{R})+(\bar{\nabla}^{2}-4R)\big{]}^{2}\Big{)}C+\frac{1}{2}(\bar{\eta}+\mbox{\boldmath$\eta$})^{2}+\frac{1}{2}(\bar{\mbox{\boldmath$\eta$}}+\eta)^{2}$
+$\phantom{=}\quad+\frac{m}{2}(\bar{\eta}^{\prime}+\eta^{\prime})^{2}-\frac{m}{8}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}^{2}\Big{\}}$
+$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\tilde{C},(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\tilde{C}\big{]}\eta$
+$\phantom{=}\quad+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\tilde{C},(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.30)
 […]
@@
 […]
-$Z$ $=\int\big{[}\mathscr{D}V\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\big{[}\mathscr{D}\bar{\eta}\big{]}\big{[}\mathscr{D}\eta\big{]}\big{[}\mathscr{D}\bar{\mbox{\boldmath$\eta$}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\eta$}\big{]}\big{[}\mathscr{D}\pi\big{]}\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)\delta(\bar{W}_{\dot{\alpha}}-\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\pi)$
-×$\qquad\qquad\times\delta(\nabla^{2}\pi)\delta(\bar{\nabla}^{2}\pi)\text{Det}(\mathcal{H})\exp\Big{(}\frac{{\rm i}}{\hbar}\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}V^{2}-2\tilde{C}\Box_{V}C+{\mbox{\boldmath$\eta$}}\bar{\eta}+{\eta}\bar{\mbox{\boldmath$\eta$}}\Big{\}}\Big{)}~{},$ (2.29)
+$Z$ $=\int\big{[}\mathscr{D}V\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\big{[}\mathscr{D}\bar{\eta}\big{]}\big{[}\mathscr{D}\eta\big{]}\big{[}\mathscr{D}\bar{\mbox{\boldmath$\eta$}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\eta$}\big{]}\big{[}\mathscr{D}\pi\big{]}\delta\Big{(}(\nabla^{2}-4\bar{R})\pi\Big{)}\delta\Big{(}(\bar{\nabla}^{2}-4R)\pi\Big{)}$
+×$\qquad\qquad\times\delta\Big{(}W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4{R})\nabla_{\alpha}\pi\Big{)}\delta\Big{(}\bar{W}_{\dot{\alpha}}-\frac{{\rm i}}{4}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\pi\Big{)}\text{Det}(\mathcal{H})$
+×$\qquad\qquad\times\exp\Big{(}{\rm i}\,\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}V^{2}-2\tilde{C}\Box_{V}C+\frac{1}{2}({\mbox{\boldmath$\eta$}}+\bar{\eta})^{2}+\frac{1}{2}({\eta}+\bar{\mbox{\boldmath$\eta$}})^{2}\Big{\}}\Big{)}~{},$ (2.31)
 where we have introduced the operators
-$\mathcal{H}$ $=\begin{pmatrix}0&-\frac{1}{4}\bar{\nabla}^{2}\\ -\frac{1}{4}\nabla^{2}&0\end{pmatrix}~{},$ (2.30a)
-$\Box_{V}$ $=-\frac{1}{16}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}-\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}$
-$=\nabla^{a}\nabla_{a}-\frac{1}{4}W^{\alpha}\nabla_{\alpha}+\frac{1}{4}\bar{W}_{\dot{\alpha}}\bar{\nabla}^{\dot{\alpha}}~{}.$ (2.30b)
+$\mathcal{H}$ $=\begin{pmatrix}R&-\frac{1}{4}(\bar{\nabla}^{2}-4R)\\ -\frac{1}{4}({\nabla}^{2}-4\bar{R})&\bar{R}\end{pmatrix}~{},$ (2.32a)
+$\Box_{V}$ $=-\frac{1}{16}\Big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}({\nabla}^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}-\big{[}({\nabla}^{2}-4\bar{R})+(\bar{\nabla}^{2}-4R)\big{]}^{2}\Big{)}$
+$=\nabla^{a}\nabla_{a}-\frac{1}{4}G^{{\alpha{\dot{\alpha}}}}[\nabla_{\alpha},\nabla_{\dot{\alpha}}]-\frac{1}{4}\bar{\nabla}_{\dot{\alpha}}\bar{R}\bar{\nabla}^{\dot{\alpha}}-\frac{1}{4}\nabla^{\alpha}R\nabla_{\alpha}-\frac{1}{4}\bar{\nabla}^{2}\bar{R}-\frac{1}{4}\nabla^{2}R$
+$\phantom{=}+2R\bar{R}-\frac{1}{4}R(\bar{\nabla}^{2}-4R)-\frac{1}{4}\bar{R}({\nabla}^{2}-4\bar{R})-\frac{1}{4}W^{\alpha}\nabla_{\alpha}+\frac{1}{4}\bar{W}_{\dot{\alpha}}\bar{\nabla}^{\dot{\alpha}}~{}.$ (2.32b)
 The operator $\mathcal{H}$ acts on the space of covariantly chiral-antichiral pairs in the adjoint representation of $G$, while $\Box_{V}$ is defined to act on the space of unconstrained scalar superfields in the adjoint representation of $G$.
+The operators $\mathcal{H}$ and $\Box_{V}$ are gauge-covariant extensions of the operators introduced in [51].
 To evaluate the $\delta$-functions in (2.3), we introduce a background quantum splitting [52] of the prepotentials in (2.6)
@@
 The field strength $W_{\alpha}$ takes the form
-$W_{\alpha}={\rm e}^{\bar{\Omega}_{Q}}\left(-\frac{1}{4}\bar{\cal D}^{2}\Big{(}{\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}\Big{)}+w_{\alpha}\right){\rm e}^{-\bar{\Omega}_{Q}}~{},$ (2.33)
+$W_{\alpha}={\rm e}^{\bar{\Omega}_{Q}}\left(-\frac{1}{4}(\bar{\cal D}^{2}-4{R})\Big{(}{\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}\Big{)}+w_{\alpha}\right){\rm e}^{-\bar{\Omega}_{Q}}~{},$ (2.35)
 where $w_{\alpha}$ is the covariantly chiral background field strength, $\bar{\cal D}_{\dot{\beta}}w_{\alpha}=0$.
 Since we are interested in evaluating the $\delta$-functions in (2.3), we decompose $W_{\alpha}$ to first order in the quantum superfields,
-$W_{\alpha}\approx-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+w_{\alpha}+\big{[}\bar{\Omega}_{Q},w_{\alpha}\big{]}~{}.$ (2.34)
+$W_{\alpha}\approx-\frac{1}{4}(\bar{\cal D}^{2}-4{R}){\cal D}_{\alpha}V_{Q}+w_{\alpha}+\big{[}\bar{\Omega}_{Q},w_{\alpha}\big{]}~{}.$ (2.36)
 The $\delta$-functions in (2.3) tell us that the background connection must be chosen to be flat,
-$w_{\alpha}=0\quad\implies\quad{\rm e}^{\bar{\omega}}={\rm e}^{\Phi}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.35)
-In evaluating $\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)$, it suffices to keep only the terms of first order in the quantum prepotentials and $\pi$,
-$\delta\left(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi\right)=\delta\left(-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+\frac{{\rm i}}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}\pi\right)~{}.$ (2.36)
+$w_{\alpha}=0\quad\implies\quad{\rm e}^{\bar{\omega}}={\rm e}^{\Phi}~{},\qquad{\bar{\mathfrak{D}}}_{\dot{\alpha}}\Phi=0~{}.$ (2.37)
+In evaluating $\delta(W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4{R})\nabla_{\alpha}\pi)$, it suffices to keep only the terms of first order in the quantum prepotentials and $\pi$,
+$\delta\left(W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\pi\right)=\delta\left(-\frac{1}{4}(\bar{{\cal D}}^{2}-4{R}){\cal D}_{\alpha}V_{Q}+\frac{{\rm i}}{4}(\bar{{\cal D}}^{2}-4{R}){\cal D}_{\alpha}\pi\right)~{}.$ (2.38)
 Now we consider the following change of variables:
-$V_{Q}$ $=\rho+\bar{\rho}+\frac{1}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}+\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\rho=0~{},$ (2.37a)
+$V_{Q}$ $=\rho+\bar{\rho}+\frac{1}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}+\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\rho=0~{},\quad{\bar{\cal D}}_{\dot{\alpha}}\lambda_{\alpha}=0~{},$ (2.39a)
 $\pi$ $=\gamma+\bar{\gamma}+\frac{{\rm i}}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}-\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\gamma=0~{}.$ (2.39b)
 Such a change of variables has a non-trivial Jacobian $\mathfrak{J}$. It can be computed by applying this change of variables in the right-hand side of
-$1=\int\big{[}\mathscr{D}V_{Q}\big{]}\big{[}\mathscr{D}\pi\big{]}\exp\Big{(}\frac{{\rm i}}{2}\text{tr}\int\,{\rm d}^{4|4}z\,\Big{\{}{V_{Q}}^{2}-\pi^{2}\Big{\}}\Big{)}$ (2.38)
+$1=\int\big{[}\mathscr{D}V_{Q}\big{]}\big{[}\mathscr{D}\pi\big{]}\exp\Big{(}\frac{{\rm i}}{2}\text{tr}\int\,{\rm d}^{4|4}z\,E\,\Big{\{}{V_{Q}}^{2}-\pi^{2}\Big{\}}\Big{)}$ (2.40)
 to result with
@@
 where ~$\widetilde{\mathscr{H}}_{c}$ (~$\widetilde{\mathscr{H}}_{a}$) is a d’Alembertian on the space of covariantly chiral (antichiral) spinor superfields
-~$\widetilde{\mathscr{H}}_{c}=\frac{1}{16}\bar{{\cal D}}^{2}{\cal D}^{2}~{},\qquad\widetilde{\mathscr{H}}_{a}=\frac{1}{16}{\cal D}^{2}\bar{{\cal D}}^{2}~{}.$ (2.40)
-Here and below, a tilde over operators, such as $\widetilde{{\cal H}}$ and ~$\widetilde{{\square}}_{V}={\cal D}^{a}{\cal D}_{a}$, indicates that these operators are constructed using the flat-connection covariant derivatives ${\cal D}_{A}$.
+~$\widetilde{\mathscr{H}}_{c}=\frac{1}{16}(\bar{{\cal D}}^{2}-4{R})({\cal D}^{2}-6\bar{R})~{},\qquad\widetilde{\mathscr{H}}_{a}=\frac{1}{16}({\cal D}^{2}-4\bar{{R}})(\bar{{\cal D}}^{2}-6{{R}})~{}.$ (2.42)
+Here and below, a tilde over operators, such as $\widetilde{{\cal H}}$, indicates that these operators are constructed using the flat-connection covariant derivatives ${\cal D}_{A}$.
 With this change of variables, we find that eq.
 (2.3) simplifies to
-$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{{\square}}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.41)
+$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{{\square}}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})\exp\Big{(}{\rm i}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.43)
 The quantum gauge field is now equal to zero, $V_{Q}=0$, and the relations (2.33) and (2.37) tell us that
-$V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.42)
-Now, repeating the analysis in [23] gives
-$\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{\square}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})=1~{},$ (2.43)
-and we obtain the in-out vacuum amplitude for the supersymmetric nonlinear $\sigma$-model (2.11)
-$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.44)
+$V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Phi=0~{}.$ (2.44)
+Now, the analysis in [23] (see also [44] for a review) implies that
+$\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{\square}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})~{},$ (2.45)
+is a topological invariant, that is it does not depend on the supergravity prepotentials $H^{m}$, $\varphi$ and $\bar{\varphi}$, where $H^{m}$ is the gravitational superfield [53, 54] and $\varphi$ the so-called chiral compensator [54] (see [44, 47] for a review).
+Therefore, the in-out vacuum amplitude (2.43) reduces to that for the supersymmetric nonlinear $\sigma$-model (2.13)
+$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\exp\Big{(}{\rm i}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.46)
 Thus we have demonstrated that the partition functions of the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.13) coincide, which implies the equality of the “on-shell” effective actions in these theories[34].
-This establishes the quantum equivalence of the two theories.
+This establishes the quantum equivalence of the two theories in the sense that the quantum supercurrents of the two theories coincide.
 One may also add a source term for $V$ to (2.1) and repeat the analysis to obtain formal equality of their partition functions.
-The analysis in this section may be generalised by coupling the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.11) to a background supergravity multiplet.
-Then (2.43) will turn into a topological invariant.
-As a result, the quantum supercurrents in the two theories $\langle J_{\alpha{\dot{\alpha}}}\rangle$ will coincide.
 This is an extension of the results obtained in [23, 42].
@@
 Here we will make use of the modified Faddeev-Popov procedure described in [23] to quantise the tensor multiplet model (3.6).
+[^10] Quantisation of this model within the BV formalism is described in appendix B.
 Our first step is to define a generalised delta-function $\hat{\delta}[H^{++}]$, with $H^{++}$ the gauge fixing function (3.12).
@@
 This equivalence may be immediately extended to the more general class of theories defined by replacing the $V^{2}$ term in (2.1) with a general function of $V$,
-$S[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\text{tr}\int{\rm d}^{4|4}z\,\mathfrak{F}(V)~{},$ (4.1)
+$S[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\text{tr}\int{\rm d}^{4|4}z\,E\,\mathfrak{F}(V)~{},$ (4.1)
 Specifically, by fixing $\mathfrak{F}(V)$ and repeating the same analysis as in section 2, one may show quantum equivalence to the following $\sigma$-model
-$S[\Phi,\bar{\Phi}]=\text{tr}\int{\rm d}^{4|4}z\,\mathfrak{F}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (4.2)
+$S[\Phi,\bar{\Phi}]=\text{tr}\int{\rm d}^{4|4}z\,E\,\mathfrak{F}({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi})~{}.$ (4.2)
 Additionally, one may instead consider the model obtained by replacing the $V^{2}$ term in (2.1) with the super Yang-Mills action
-$S_{\text{T}}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\Big{\{}\chi^{\alpha}W_{\alpha}+\frac{1}{g^{2}}W^{2}\Big{\}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{g^{2}}\bar{W}^{2}\Big{\}}~{}.$ (4.3)
+$S_{\text{T}}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\Big{\{}\chi^{\alpha}W_{\alpha}+\frac{1}{g^{2}}W^{2}\Big{\}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\Big{\{}\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{g^{2}}\bar{W}^{2}\Big{\}}~{}.$ (4.3)
 In addition to invariance under (2.7), this model enjoys super Yang-Mills gauge symmetry [^11] We recall that $L_{A}B=[A,B]$, for operators $A$ and $B$.
-$\delta_{\Lambda}V=-{{\rm i}}L_{V}(\Lambda+\bar{\Lambda})+{{\rm i}}L_{V}\coth(L_{V})(\bar{\Lambda}-\Lambda)~{},\qquad\bar{D}_{\dot{\alpha}}\Lambda=0~{}.$ (4.4)
+$\delta_{\Lambda}V=-{{\rm i}}L_{V}(\Lambda+\bar{\Lambda})+{{\rm i}}L_{V}\coth(L_{V})(\bar{\Lambda}-\Lambda)~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Lambda=0~{}.$ (4.4)
 Further, it is classically topological; it describes trivial dynamics.
 This property proves to also hold at the quantum level, which may be seen via a slight modification of the analysis of section 2.
-Specifically, it is necessary to incorporate the ghost action corresponding to (4.4) and choose a gauge fermion $\Psi(\mbox{\boldmath$\Phi$})$ such that the gauge $D^{2}V=\bar{D}^{2}V=0$ is enforced.
+Specifically, it is necessary to incorporate the ghost action corresponding to (4.4) and choose a gauge fermion $\Psi(\mbox{\boldmath$\Phi$})$ such that the gauge $(\mathfrak{D}^{2}-4\bar{R})V=(\bar{\mathfrak{D}}^{2}-4R)V=0$ is enforced.
 With these adjustments, a similar analysis to that conducted in section 2 leads to $Z=1$ for this theory, implying trivial dynamics.
@@
 $\bar{\cal D}^{\dot{\alpha}}_{i}\Psi=0\quad\implies\quad\bar{\Delta}\Delta\Psi=\Big{\{}({\cal D}^{a}{\cal D}_{a})^{2}+\dots\Big{\}}\Psi~{}.$ (4.8)
+An ${\cal N}=2$ locally supersymmetric analogue of the topological model (4.3) is
+$S=\frac{1}{2}\int{\rm d}^{4}x{\rm d}^{4}\theta\,{\cal E}\,W\big{(}\Psi+\frac{1}{g^{2}}W\big{)}+{\rm c.c.}~{},$ (4.9)
+where the prepotential $\Psi$ for the tensor multiplet is now covariantly chiral, and chiral field strength $W$ for the vector multiplet obeys the constraints
+${\bar{\cal D}}^{\dot{\alpha}}_{i}W=0~{},\qquad\Big{(}{\cal D}^{\alpha(i}{\cal D}_{\alpha}^{j)}+4S^{ij}\Big{)}W=\Big{(}{\bar{\cal D}}_{\dot{\alpha}}{}^{(i}{\bar{\cal D}}^{j){\dot{\alpha}}}+4\bar{S}^{ij}\Big{)}\bar{W}~{}.$ (4.10)
+This model respects the gauge symmetries of both tensor and vector multiplets.
+It would be interesting to compute the partition function for the model since its logarithmically divergent part should contain the ${\cal N}=2$ Gauss-Bonnet term [74].
 In the supersymmetric literature, the BV procedure has been employed [89, 90, 91, 92] to quantise theories involving the massless non-minimal scalar multiplet [93] described by a complex linear superfield.
@@
 $D^{\pm}_{\alpha}=D^{i}_{\alpha}\,u^{\pm}_{i}~{},\qquad{\bar{D}}^{\pm}_{\dot{\alpha}}={\bar{D}}^{i}_{\dot{\alpha}}\,u^{\pm}_{i}~{}.$ (A.3)
-The operators $D^{+}_{\alpha}$ and $\bar{D}^{+}_{\dot{\alpha}}$ strictly anticommute, and therefore on can introduce analytic superfields of $\mathsf{U}(1)$ charge $n$, $\varphi^{(n)}(z,u)$, with the properties:
+The operators $D^{+}_{\alpha}$ and $\bar{D}^{+}_{\dot{\alpha}}$ strictly anticommute, and therefore one can introduce analytic superfields of $\mathsf{U}(1)$ charge $n$, $\varphi^{(n)}(z,u)$, with the properties:
 $D^{+}_{\alpha}\varphi^{(n)}=0~{},\qquad\bar{D}^{+}_{\dot{\alpha}}\varphi^{(n)}=0;$ (A.4a)
@@
 […]
+Appendix B BV quantisation of ${\cal N}=2$ gauge theories
+This appendix is devoted to the BV quantisation of two ${\cal N}=2$ supersymmetric gauge theories in harmonic superspace.
+Specifically, we will consider the Abelian vector [^12] Quantisation of ${\cal N}=2$ super Yang-Mills theory was performed in [72] in the framework of the background field method.
+For an Abelian gauge group, its results are in agreement with the ones presented below.
+and tensor multiplets.
+While the latter was quantised in section 3 via a modified Faddeev-Popov procedure, we hope that this technical appendix is of use to the reader unfamiliar with the formalism.
+B.1 The vector multiplet
+The ${\cal N}=2$ vector multiplet is known to be described by a real analytic superfield $V^{++}$
+$D^{+}_{\alpha}V^{++}=0~{},\qquad\bar{D}^{+}_{\dot{\alpha}}V^{++}=0~{},$ (B.1)
+which is defined modulo the following Abelian gauge transformations
+$\delta_{\lambda}V^{++}=-D^{++}\lambda~{},\qquad D^{+}_{\alpha}\lambda=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}\lambda=0~{}.$ (B.2)
+It may then be shown that the following action for $V^{++}$
+$S_{\text{VM}}[V^{++}]=\frac{1}{2}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{V^{++}(1)V^{++}(2)}{(u_{1}^{+}u_{2}^{+})^{2}}~{},$ (B.3)
+is inert under (B.2). Below, we will quantise this model by employing the BV formalism.
+In accordance with eq. (2.23), the minimal solution to the master equation (2.20) is
+$S[\mbox{\boldmath$\Phi$}_{\text{min}},\mbox{\boldmath$\Phi$}^{*}_{\text{min}}]=S_{\text{VM}}[V^{++}]+\int{\rm d}\zeta^{(-4)}\,V^{*++}D^{++}C~{},\qquad\mbox{\boldmath$\Phi$}_{\text{min}}=\{V^{++},C\}~{},$ (B.4)
+[…]
+$(\tilde{C},\pi)~{},\quad D_{\alpha}^{+}C=D_{\alpha}^{+}\pi=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}C=\bar{D}_{\dot{\alpha}}^{+}\pi=0~{},\quad\epsilon(\tilde{C})=\epsilon(\pi)+1=1~{}.$ (B.5)
+Then, the following functional [^13] Given an analytic superfield of charge $n$, its corresponding antifield carries charge $4-n$.
+$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]=S[\mbox{\boldmath$\Phi$}_{\text{min}},\mbox{\boldmath$\Phi$}^{*}_{\text{min}}]+\int{\rm d}\zeta^{(-4)}\,\tilde{C}^{*(4)}\pi~{},$ (B.6)
+is also a solution of the master equation and satisfies $\Delta S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]=0$. Now, we choose the gauge fermion to be
+$\Psi[\mbox{\boldmath$\Phi$};\Lambda^{(4)}]=\int{\rm d}\zeta^{(-4)}\,\Big{(}\tilde{C}D^{++}V^{++}-\tilde{C}\Lambda^{(4)}\Big{)}~{},\qquad D_{\alpha}^{+}\Lambda^{(4)}=0~{},\quad\bar{D}_{\dot{\alpha}}^{+}\Lambda^{(4)}=0~{},$ (B.7)
+where $\Lambda^{(4)}$ is a background analytic superfield. Eliminating antifields in (B.6) in accordance with (B.7) yields the following
+$S\Big{[}\mbox{\boldmath$\Phi$},\frac{\delta\Psi}{\delta\mbox{\boldmath$\Phi$}}\Big{]}=S_{\text{VM}}[V^{++}]+\int{\rm d}\zeta^{(-4)}\,(D^{++}V^{++}-\Lambda^{(4)})\pi~{}.$ (B.8)
+This analysis indicates that the in-out vacuum amplitude for (B.3) is
+$Z=\int\big{[}\mathscr{D}V^{++}\big{]}\,\delta(D^{++}V^{++}-\Lambda^{(4)})\exp\Big{(}{\rm i}S_{\text{VM}}[V^{++}]\Big{)}~{}.$ (B.9)
+Since the right-hand side is independent of the analytic superfield $\Lambda^{(4)}$, we can integrate over it with the weight
+$\tilde{\Delta}_{\rm NK}\exp\left\{\frac{{\rm i}}{2\alpha}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Lambda^{(4)}(1)\,\frac{(u^{-}_{1}u^{-}_{2})}{(u^{+}_{1}u^{+}_{2})^{3}}\,\Lambda^{(4)}(2)\right\}~{},\qquad\alpha\in\mathbb{R}-\{0\}~{},$ (B.10)
+where $\tilde{\Delta}_{\rm NK}$ is a Nielsen-Kallosh determinant. The latter is determined by requiring
+$1=\tilde{\Delta}_{\rm NK}\int\big{[}\mathscr{D}\Lambda^{(4)}\big{]}\,\exp\left\{{{\rm i}}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Lambda^{(4)}(1)\,\frac{(u^{-}_{1}u^{-}_{2})}{(u^{+}_{1}u^{+}_{2})^{3}}\,\Lambda^{(4)}(2)\right\}~{}.$ (B.11)
+In accordance with the analysis of [72], it is given by
+$\tilde{\Delta}_{\rm NK}=\Big{(}\text{Det}[(D^{++})^{2}]\Big{)}^{-\frac{1}{2}}\,(\text{Det}_{(0,4)}\Box)^{\frac{1}{2}}~{}.$ (B.12)
+A routine manipulation of the resulting path integral leads to
+$Z$ $=\int\big{[}\mathscr{D}V^{++}\big{]}(\text{Det}[(D^{++})^{2}])^{\frac{1}{2}}\,(\text{Det}_{(0,4)}\Box)^{\frac{1}{2}}$
+×$\qquad\qquad\times\exp\Big{(}{\rm i}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{1}{2}\Big{(}1+\frac{1}{\alpha}\Big{)}\frac{V^{++}(1)V^{++}(2)}{(u_{1}^{+}u_{2}^{+})^{2}}-\frac{{\rm i}}{2\alpha}\int{\rm d}\zeta^{(-4)}\,V^{++}\Box V^{++}\Big{]}\Big{)}~{}.$ (B.13)
+Finally, fixing $\alpha=-1$ yields Feynman gauge and we obtain
+$Z=\Big{(}\text{Det}[(D^{++})^{2}]\Big{)}^{\frac{1}{2}}\,(\text{Det}_{(0,4)}\Box)^{\frac{1}{2}}\,(\text{Det}_{(2,2)}\Box)^{-\frac{1}{2}}~{}.$ (B.14)
+B.2 The tensor multiplet
+We recall that the ${\cal N}=2$ tensor multiplet is described by action (3.6), which we reiterate for convenience
+$S_{\text{TM}}[\Psi,\bar{\Psi}]=\frac{1}{2}\int{\rm d}\zeta^{(-4)}\,(G^{++})^{2}~{},\qquad G^{++}(z,u)=\frac{1}{8}(D^{+})^{2}\Psi(z)+\frac{1}{8}({\bar{D}}^{+})^{2}{\bar{\Psi}}(z)~{},$ (B.15)
+where $\Psi$ is a chiral superfield; $\bar{D}_{\dot{\alpha}}^{i}\Psi=0$.
+This action is invariant under the Abelian gauge transformations (3.3), which can be expressed in terms of a real analytic superfield (3.18)
+$\delta_{\mathfrak{U}}\Psi=\frac{{\rm i}}{4}\int{\rm d}u\,({\bar{D}}^{-})^{2}{\mathfrak{U}}^{++}~{},\qquad D^{+}_{\alpha}{\mathfrak{U}}^{++}=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}{\mathfrak{U}}^{++}=0~{}.$ (B.16)
+This gauge symmetry is reducible as (B.16) is inert under
+$\delta_{\sigma}{\mathfrak{U}}^{++}=D^{++}\sigma~{},\qquad D^{+}_{\alpha}\sigma=0~{},\quad\bar{D}^{+}_{\dot{\alpha}}\sigma=0~{}.$ (B.17)
+Below, we will employ the BV scheme to quantise this model and provide an alternative proof of quantum equivalence to the $\omega$ hypermultiplet model (3.16) to the one given in section 3.3.
+Making use of (B.16) and (B.17) in conjunction with (2.23), we obtain the minimal solution to the master equation (2.20)
+$S[\mbox{\boldmath$\Phi$}_{\text{min}},\mbox{\boldmath$\Phi$}^{*}_{\text{min}}]$ $=S_{\text{TM}}[\Psi,\bar{\Psi}]+\frac{{\rm i}}{4}\bigg{\{}\int{\rm d}^{4|4}z_{+}{\rm d}u\,\Psi^{*}(\bar{D}^{-})^{2}C^{++}-\int{\rm d}^{4|4}z_{-}{\rm d}u\,\bar{\Psi}^{*}(D^{-})^{2}C^{++}\bigg{\}}$
+$\quad+\int{\rm d}\zeta^{(-4)}\,C^{*++}D^{++}\eta~{},$ (B.18)
+[…]
+$(i):$ $\qquad(\tilde{C}^{++},\pi^{++})~{},\quad\epsilon(\tilde{C}^{++})=\epsilon(\pi^{++})+1=1~{},$ (B.19a)
+$(ii):$ $\qquad(\tilde{\eta},\tilde{\pi})~{},\quad\epsilon(\tilde{\eta})=\epsilon(\tilde{\pi})-1=0~{},$ (B.19b)
+$(iii):$ $\qquad({\eta}^{\prime},{\pi}^{\prime})~{},\quad\epsilon({\eta}^{\prime})=\epsilon({\pi}^{\prime})-1=0~{}.$ (B.19c)
+Then, the following functional
+$S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]=S[\mbox{\boldmath$\Phi$}_{\text{min}},\mbox{\boldmath$\Phi$}^{*}_{\text{min}}]+\int{\rm d}\zeta^{(-4)}\,\Big{(}\tilde{C}^{*++}\pi^{++}+\tilde{\eta}^{*(4)}\tilde{\pi}+\eta^{\prime*(4)}\pi^{\prime}\Big{)}~{},$ (B.20)
+is also a solution of the master equation and satisfies $\Delta S[\mbox{\boldmath$\Phi$},\mbox{\boldmath$\Phi$}^{*}]=0$.
+We choose the gauge fermion to be [^15] Here we denote the gauge fermion by $\Psi_{\text{GF}}$ to distinguish it from the tensor multiplet’s prepotential.
+$\Psi_{\text{GF}}[\mbox{\boldmath$\Phi$};\Omega^{(4)},\Upsilon^{(4)}]$ $=\int{\rm d}\zeta^{(-4)}\,\Big{(}\tilde{C}^{++}H^{++}-\tilde{\eta}D^{++}C^{++}+D^{++}\tilde{C}\eta^{\prime}$
+$\qquad\qquad\qquad\qquad\qquad-\frac{1}{2m}\tilde{C}^{++}\pi^{++}+\Omega^{(4)}\tilde{\eta}-\Upsilon^{(4)}\eta^{\prime}\Big{)}~{},$ (B.21)
+where $H^{++}$ is given by (3.12) while $\Omega^{(4)}$ and $\Upsilon^{(4)}$ are fermionic background analytic superfields.
+Eliminating antifields in (B.20) in accordance with (B.2) yields the following
+$S\Big{[}\mbox{\boldmath$\Phi$},\frac{\delta\Psi_{\text{GF}}}{\delta\mbox{\boldmath$\Phi$}}\Big{]}$ $=S_{\text{TM}}[\Psi,\bar{\Psi}]-\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{\tilde{C}^{++}(1)C^{++}(2)}{(u_{1}^{+}u_{2}^{+})^{2}}+\int{\rm d}\zeta^{(-4)}\,\Big{(}D^{++}\tilde{\eta}D^{++}\eta$
+$+\Big{(}H^{++}-D^{++}\eta^{\prime}-\frac{1}{2m}\pi^{++}\Big{)}\pi^{++}+\Big{(}\Omega^{(4)}-D^{++}C^{++}\Big{)}\tilde{\pi}+\Big{(}D^{++}\tilde{C}^{++}-\Upsilon^{(4)}\Big{)}\pi^{\prime}\Big{)}~{}.$ (B.22)
+This analysis indicates that the in-out vacuum amplitude for (B.3) is
+$Z$ $=\int\big{[}\mathscr{D}\bar{\Psi}\big{]}\big{[}\mathscr{D}{\Psi}\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}{C}\big{]}\big{[}\mathscr{D}\tilde{\eta}\big{]}\big{[}\mathscr{D}{\eta}\big{]}\big{[}\mathscr{D}\eta^{\prime}\big{]}\,\delta(D^{++}\tilde{C}^{++}-\Omega^{(4)})\delta(D^{++}{C}^{++}-\Upsilon^{(4)})$
+×$\qquad\times\exp\Big{(}{\rm i}\Big{\{}S_{\text{TM}}[\Psi,\bar{\Psi}]-\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\frac{\tilde{C}^{++}(1)C^{++}(2)}{(u_{1}^{+}u_{2}^{+})^{2}}+\int{\rm d}\zeta^{(-4)}\,\Big{[}-\tilde{\eta}(D^{++})^{2}\eta$
+$\qquad\qquad\qquad-\frac{m}{2}\eta^{\prime}(D^{++})^{2}\eta^{\prime}+\frac{m}{2}(H^{++})^{2}\Big{]}\Big{\}}\Big{)}~{},$ (B.23)
+where we have integrated out $\pi^{++}$, $\tilde{\pi}$ and $\pi^{\prime}$.
+Since the right-hand side is independent of the background analytic superfields $\Omega^{(4)}$ and $\Upsilon^{(4)}$, we can integrate over them with the weight
+$({\Delta}_{\rm NK})^{-1}\exp\left\{\frac{{\rm i}}{n}\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Omega^{(4)}(1)\,\frac{(u^{-}_{1}u^{-}_{2})}{(u^{+}_{1}u^{+}_{2})^{3}}\,\Upsilon^{(4)}(2)\right\}~{},\qquad n\in\mathbb{R}-\{0\}~{},$ (B.24)
+where ${\Delta}_{\rm NK}$ is the Nielsen-Kallosh determinant (3.38).
+[^16] It should be noted that the background fields in (B.24) are of opposite statistics to those of (3.36).
+This is why their Nielsen-Kallosh determinants are inverse to each other.
+The resulting path integral is
+$Z$ $=\int\big{[}\mathscr{D}\bar{\Psi}\big{]}\big{[}\mathscr{D}{\Psi}\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}{C}\big{]}\big{[}\mathscr{D}\tilde{\eta}\big{]}\big{[}\mathscr{D}{\eta}\big{]}\big{[}\mathscr{D}\eta^{\prime}\big{]}\,({\Delta}_{\rm NK})^{-1}\exp\Big{(}{\rm i}\Big{\{}S_{\text{TM}}[\Psi,\bar{\Psi}]$
+$\qquad+\int{\rm d}^{4|8}z{\rm d}u_{1}{\rm d}u_{2}\,\Big{(}\frac{1}{n}-1\Big{)}\frac{\tilde{C}^{++}(1)C^{++}(2)}{(u_{1}^{+}u_{2}^{+})^{2}}+\frac{1}{2n}\int{\rm d}\zeta^{(-4)}\,\tilde{C}^{++}\Box C^{++}$
+$\qquad\qquad+\int{\rm d}\zeta^{(-4)}\,\Big{[}-\tilde{\eta}(D^{++})^{2}\eta-\frac{m}{2}\eta^{\prime}(D^{++})^{2}\eta^{\prime}+\frac{m}{2}(H^{++})^{2}\Big{]}\Big{\}}\Big{)}~{}.$ (B.25)
+Choosing the gauge $m=n=1$ leads to significant simplifications in (B.2) and allows one to readily perform the path integral over all remaining fields. This leads to
+$Z$ $=({\Delta}_{\rm NK})^{-1}\left({\rm Det}\,{\cal H}\right)^{-1/2}({\rm Det}_{(2,2)}\,{\square}{})\,\Big{(}{\rm Det}\big{[}(D^{++})^{2}\big{]}\Big{)}^{-3/2}$
+$=({\rm Det}\,{\cal H})^{-1/2}({\rm Det}_{(2,2)}\,{\square}{})\,({\rm Det}_{(0,4)}\,{\square}{})^{-1}\,\Big{(}{\rm Det}\big{[}(D^{++})^{2}\big{]}\Big{)}^{-1/2}~{},$
+$=\Big{(}\text{Det}[(D^{++})^{2}]\Big{)}^{-\frac{1}{2}}~{},$ (B.26)
+where we have made use of eq. (3.48) in the final line. This result corroborates the quantum equivalence of the tensor and $\omega$ hypermultiplet demonstrated in section 3.3.